### v1

# A Survey on Graph Neural Network Acceleration: Algorithms, Systems, and Customized Hardware

Abstract.
Graph neural networks (GNNs) are emerging for machine learning research on graph-structured data. GNNs achieve state-of-the-art performance on many tasks, but they face scalability challenges when it comes to real-world applications that have numerous data and strict latency requirements. Many studies have been conducted on how to accelerate GNNs in an effort to address these challenges. These acceleration techniques touch on various aspects of the GNN pipeline, from smart training and inference algorithms to efficient systems and customized hardware. As the amount of research on GNN acceleration has grown rapidly, there lacks a systematic treatment to provide a unified view and address the complexity of relevant works. In this survey, we provide a taxonomy of GNN acceleration, review the existing approaches, and suggest future research directions. Our taxonomic treatment of GNN acceleration connects the existing works and sets the stage for further development in this area.
graph neural networks, model acceleration, hardware acceleration/accelerator
copyright: acmcopyrightjournalyear: XXXXdoi: XXXXXXX.XXXXXXXjournal: CSURccs: Computing methodologies Neural networksccs: Computer systems organization Distributed architecturesccs: Computer systems organization Data flow architecturesccs: Hardware Integrated circuits
1. Introduction
Graphs are a natural and powerful data structure for representing entities and their relationships. Many real-world data can be represented as graphs with nodes denoting a collection of entities and edges denoting their pairwise relationships, such as individuals in social networks, financial transactions between firms and banks, atoms and bonds in molecules, and vehicles in transportation systems. Graph neural networks (GNNs) (Kipf and Welling, 2016; Veličković et al., 2017; Hamilton et al., 2017a) have recently become the most widely used graph machine learning (ML) model for learning knowledge and making predictions on graph data. GNNs have achieved state-of-the-art performance in many graph ML applications. They are used, for example, in recommendations on social graphs (Zhang et al., 2019a; Mao et al., 2021; Wu et al., 2020b), fraud account detection on financial graphs (Dou et al., 2020), drug discoveries from molecule graphs (Jiang et al., 2021), traffic forecasting on transportation graphs (Jiang and Luo, 2022), and so on.
The superior performance of GNNs on graphs is mainly due to their ability to combine the entity information, represented as the node features, and the relationships, represented as the graph structure. GNNs take node features and the graph structure as input and output node representations, which can be used for various graph ML tasks like node classification, edge prediction, and graph classification. For each node in a graph, a GNN learns its representation by message passing, which aggregates messages (features) from the node’s neighbors and updates the representation by transforming the aggregated messages with a neural network. Each aggregate-and-update operation is called a GNN layer, which allows the node representation to combine information from the node’s direct neighbors. Stacking multiple GNN layers will apply such operations recursively so the node representation can gain information from multi-hop neighbors. Thus, the representation can capture the contextual information of a local neighborhood and be used to answer complex ML questions on graphs. For a target node with an $L$-layer GNN applied on it, all of its neighbors within $L$ hops form the receptive field of the target node. A recursively constructed $L$-layer tree graph with the target node as the root and the neighbors of each node as its children is called a computation graph of the target node.
As the field of graph ML quickly develops, graph data is becoming gigantic with numerous nodes and edges in a single graph. For example, the Twitter user graph has 288M monthly active users (nodes) and an estimated 208 follow relations (edges) per user as of 3/2015, and the Facebook user graph has 1.39B active users and more than 400B total edges as of 12/2014 (Ching et al., 2015). These numbers have grown much larger and become hard to estimate. The commonly used academic benchmarks have also increased from thousands of nodes, e.g., Cora (Sen et al., 2008a), to as many as 240M nodes in the recent Open Graph Benchmark (OGB) (Hu et al., 2021). Large graphs raise scalability challenges for both GNN training and inference. Although deep neural networks (DNNs) for other data types also face scalability challenges when the number of DNN parameters becomes large, the challenge for GNNs is unique and there is greater emphasis on the large graph data (Zhang et al., 2020a; Jia et al., 2020b) rather than the large number of model parameters. When training DNNs on images or text, even if the amount of data is large, data can be processed by randomly sampling mini-batches because all data instances are assumed to be independent and identically distributed (iid). The DNN inference can also be done independently for each instance. For GNNs on graphs, however, the nodes are dependent on each other, which brings both benefits and challenges. GNNs leverage node dependency to learn informative representations and make predictions, but node dependency also means multi-hop neighbors are involved in the GNN computation for every single node. Therefore, it is nontrivial to partition graph nodes into mini-batches for GNN training considering the node dependency. Moreover, node dependency means that the sizes of computation graphs grow exponentially in the number of GNN layers (Alon and Yahav, 2021). On large graphs where deeper GNNs are often needed for better performance (Chen et al., 2020a; Li et al., 2021), not only does the mini-batch training need to process a huge computation graph but so does the inference on a single node. Take the Twitter user graph as an example, with 208 neighbors per node, a computation graph of a three-layer GNN involves all neighbors of a node within three steps, which can contain millions of nodes (i.e., roughly $208^{3}$ minus overlapped nodes). This node dependency problem makes applying GNNs on large graphs very challenging. A different situation for large datasets is a graph database with many graphs, for example, chemical molecules. Since graphs in these databases often only have dozens to hundreds nodes, the node dependency problem is not a big concern, and they are not the focus of this survey.
Scalability challenges also show up for GNN systems, including both Commercial-off-the-Shelf (COTS) systems and customized hardware. Large computation graphs need to be processed because of the node dependency problem, but memory consumption becomes infeasible when the graph structure is represented as a huge adjacency matrix. In reality, the graph structure often needs to be processed in the sparse matrix format. However, sparse matrices have different computation patterns from standard DNNs. Therefore, the driving force of the success of deep learning on large data, i.e., the parallelism of computation devices like GPUs, cannot be fully exploited for efficient computation. This demands efficient systems with accelerated computation kernels and customized hardware accelerators like FPGA.
The scalability challenges of GNNs require acceleration techniques that will lead to resource-efficient GNN training and fast GNN inference for latency-constraint applications. Acceleration techniques can also benefit useful applications like neural architecture search. Many GNN acceleration techniques with different emphases have been proposed in the literature, including smart training and inference algorithms, efficient systems, and customized hardware designs. In this survey, we provide a unified view of different GNN acceleration techniques to address their complexity. We show a taxonomy of GNN acceleration techniques in Figure 1 that categorizes GNN acceleration techniques into three categories: (1) algorithms, (2) COTS systems, and (3) customized hardware. The algorithms category includes training acceleration techniques that modify the graph or sample from the graph, alleviating the node dependency problem. The category also includes inference acceleration techniques like pruning, quantization, and distillation, which can convert a trained GNN into a simpler and faster model. The system category includes GPU kernel acceleration for sparse matrix operations, user-defined function optimization to generate code better adjusted for the target hardware, and the design of scalable systems for distributed training on multiple devices. The customized hardware category includes hardware accelerators with different properties that cover both cases of allowing or forbidding layer customization, and have different levels of parallelization and sparsity support. Besides the three main categories, we also discuss techniques that accelerate GNNs on special heterogeneous and dynamic graphs. We discuss all the techniques in detail with regard to their advantages, applicability, and limitations. We also suggest research directions to set the stage for further development of this research topic.
Figure 1. A taxonomy of GNN acceleration. Specifically, we discuss training algorithms in Section 3, inference algorithms in Section 4, COTS systems in Section 5, customized hardware in Section 6, and special graphs and GNNs in Section 7.
2. Notations and Preliminaries
In this section, we clarify the common notations used throughout the paper in Table 1. Also, we review graph machine learning with GNNs by going through the graph data representation, the GNN model architecture, and the machine learning tasks GNNs can solve.
Table 1. Major Notations
Category Notations Descriptions
Graphs ${\mathcal{G}}$ A graph with feature-enriched nodes
$N$ The total number of nodes in ${\mathcal{G}}$, i.e., $|{\mathcal{V}}|$
${\mathcal{V}}$ The node set of ${\mathcal{G}}$
${\mathcal{E}}$ The edge set of ${\mathcal{G}}$
${\bm{X}}$ Raw node features in ${\mathcal{R}}^{N\times D}$
${\bm{x}}_{i}$ The $i$th row of ${\bm{X}}$, which is the raw feature of the node $i\in{\mathcal{V}}$
${\bm{A}}$ The adjacency matrix. ${\bm{A}}\in\{0,1\}^{N\times N}$. ${\bm{A}}_{i,j}=1$ if edge $(i,j)\in{\mathcal{E}}$, otherwise 0
$\tilde{\bm{A}}$ The adjacency matrix with self-loops. $\tilde{\bm{A}}={\bm{A}}+{\bm{I}}$
${\bm{D}}$, $\tilde{\bm{D}}$ Degree matrices of ${\bm{A}}$ and $\tilde{\bm{A}}$. ${\bm{D}}_{i,i}=\sum_{j}{\bm{A}}_{j,i}$ and ${\bm{D}}_{i,j}=0$ if $i\neq j$. Similarly for $\tilde{\bm{D}}$
${\mathcal{N}}(v)$ Neighbors of a node $v\in{\mathcal{V}}$. ${\mathcal{N}}(v)=\{u\in{\mathcal{V}}|(u,v)\in{\mathcal{E}}\}$
$d(v)$ The degree of node $v$. $d(v)=|{\mathcal{N}}(v)|$
GNNs $\textsc{GNN}_{\theta}$ The GNN model with parameters $\theta$
$L$ Total number of layers of a GNN
${\bm{H}}^{(l)}$ Hidden node representations at layer $l$. ${\bm{H}}^{(0)}={\bm{X}}$. ${\bm{H}}^{(l)}\in{\mathcal{R}}^{N\times F}$ (assume $F$ is shared for $l\geq 1$)
${\bm{h}}_{i}^{(l)}$ The $i$th row of ${\bm{H}}^{(l)}$, which is the hidden feature of node $i$ at layer $l$
$\phi(\cdot,\cdot)$ The message (a.k.a. feature extraction) function which constructs a message between two nodes
${\bm{m}}_{u,v}^{(l)}$ The constructed message from node $u$ to node $v$ at layer $l$
AGGR The aggregation function, usually order invariant, e.g., mean
${\bm{a}}_{v}^{(l)}$ The aggregated feature of node $v$ at layer $l$
UPDATE The update function. A neural network. Usually a multi-layer perceptron (MLP)
${\bm{W}}^{(l)}$ Neural network weights need to be learned in the UPDATE function at layer $l$
$\sigma$ The nonlinear activation function of the UPDATE neural network
${\bm{P}}$ The filter matrix for aggregating messages, usually a function of ${\bm{A}}$ and ${\bm{D}}$
Labels & Training ${\bm{Y}}$ One-hot node labels to predict. ${\bm{Y}}\in\{0,1\}^{N\times C}$
${\bm{y}}_{i}$ The $i$th row of ${\bm{Y}}$, which is the one-hot node label for the node $i$
$\hat{\bm{y}}_{i}$ Prediction of the node $i\in{\mathcal{V}}$, a probability vector. $\hat{\bm{y}}_{i}\in[0,1]^{C}$
ONEHOT One-hot vector function. Takes $c\in\{1,\dots,C\}$ and outputs ${\bm{y}}\in\{0,1\}^{C}$ with only ${\bm{y}}_{c}=1$
${\mathcal{L}}({\bm{y}}_{i},\hat{\bm{y}}_{i})$ The loss function for training $\textsc{GNN}_{\theta}$
2.1. Graph Data Representation
An attributed graph ${\mathcal{G}}$ includes the graph structure $({\mathcal{V}},{\mathcal{E}})$ and node features ${\bm{X}}$, with ${\mathcal{V}}$ denotes a set of nodes, and ${\mathcal{E}}$ denotes a set of edges. A straightforward way to represent the graph structure (edges ${\mathcal{E}}$) is to use an adjacency matrix ${\bm{A}}\in{\mathcal{R}}^{|{\mathcal{V}}|\times|{\mathcal{V}}|}$, with $|{\mathcal{V}}|$ denotes the number of nodes. Many graph algorithms, including the matrix version of GNNs, are presented in terms of ${\bm{A}}$. However, naively representing edges with the full matrix ${\bm{A}}$ requires $O(|{\mathcal{V}}|^{2})$ space complexity, which is impractical to implement for large graphs. In practice, most of the graphs are sparse, i.e., $|{\mathcal{E}}|<<|{\mathcal{V}}|^{2}$. Leveraging the sparsity to represent ${\mathcal{G}}$ in more efficient formats can help save memory and accelerate many computations. One popular sparse representation is the Coordinate (COO) (Wang et al., 2019b) format, which is also known as the “triplet” format. For this format, each edge $(i,j)\in{\mathcal{E}}$ is represented as a triplet $(i,j,w_{ij})$ with $w_{ij}$ being the edge weight for weighted graphs, and $w_{ij}$ equals one for unweighted graphs. The space complexity of a COO graph scales only with the number of edges, e.g., $O(|{\mathcal{E}}|)$. Another popular choice is the Compressed Sparse Row (CSR) format (Saad, 2003; Wang et al., 2019b). It is similar to the COO format but compresses those $|{\mathcal{E}}|$ triplets into three $|{\mathcal{E}}|$-dimensional arrays. One array for all the row indices, e.g., the first entry in COO. One array for all the column indices, e.g., the second entry in COO. One array for all the edge weights, e.g., the third entry in COO. The space complexity of a CSR graph is also $O(|{\mathcal{E}}|)$. Additionally, graphs often come with node features, which are represented as a feature matrix ${\bm{X}}\in{\mathcal{R}}^{|{\mathcal{V}}|\times D}$. ${\bm{X}}$ is often referred to as the raw features to distinguish them from the hidden features learned by the model. Similarly, there could be features for graph edges, but we do not assume edge features are given in this survey because they are not generally available.
2.2. Graph Neural Networks
GNN is a family of neural network models that achieve state-of-the-art performance on many machine learning tasks on graphs. GNNs take a graph and node features as input and output node representations. The node representations can be used for various downstream ML tasks like node classification and edge prediction. Node representations can also be aggregated over the whole graph to get a graph representation for graph classification.
There are many different GNN model architectures. For example, the graph convolutional network (GCN) (Kipf and Welling, 2016), graph information network (GIN) (Xu et al., 2018a), graph attention network (GAT) (Veličković et al., 2017), etc. Most of the model variants belong to the message-passing GNN framework (Gilmer et al., 2017). Message passing is a mechanism that iteratively updates the hidden representation of a node by aggregating messages from its neighbors and transforming the aggregated messages with a neural network. Each message-passing iteration is called a GNN layer, which allows the hidden representation to combine information from the node’s direct neighbors. Stacking multiple GNN layers allows multi-hop neighbors to pass messages to the central node. Thus, the node representation can capture the contextual information of a local neighborhood and be used to answer complex ML questions on graphs.
We now describe the detailed message-passing operation using the notations defined in Table 1. For layer $l$, the hidden representation ${\bm{h}}^{(l)}$ of a node $v$ is updated by first aggregating messages (a function of hidden representations from the previous iteration) from its neighbors and then combining the aggregated messages with its own hidden representation. The message-passing operation:
(1) ${\bm{h}}_{v}^{(l)}=\textsc{UPDATE}({\bm{h}}_{v}^{(l-1)},\textsc{AGGR}(\{\phi({\bm{h}}_{v}^{(l-1)},{\bm{h}}_{u}^{(l-1)})|u\in{\mathcal{N}}(v)\})).$
where $\phi(\cdot,\cdot)$ constructs the message between two nodes from their hidden representations; AGGR is the aggregation function, usually being order-invariant, e.g., mean; and UPDATE is the update function, usually being a multi-layer perceptron (MLP) with learnable weights ${\bm{W}}^{(l)}$. Equation (1) shows one message passing step for layer $l$, i.e., the $l$-th GNN layer. For a GNN with $L$ total layers, message passing is applied $L$ times to aggregate messages from neighbors within $L$-hops of $v$ to compute the representation ${\bm{h}}^{(L)}_{v}$. All the neighbors within $L$-hop of $v$ form the receptive field of $v$. A recursively constructed $L$-layer tree graph with $v$ as the root and the neighbors of each node as its children are called a computation graph of $v$.
For many GNNs, the message-passing operation in Equation (1) can be implemented by a clean matrix multiplication to update hidden representations of all $v\in{\mathcal{V}}$ at once, which is also called the full-batch version. The full-batch updates:
(2) ${\bm{H}}^{(l)}$ $=\sigma({\bm{P}}{\bm{H}}^{(l-1)}{\bm{W}}^{(l)}),$
where the matrix ${\bm{H}}^{(l-1)}$ contains hidden representations of all nodes at the $(l-1)$-th step. It is first multiplied by a filter matrix ${\bm{P}}$ (on the left) and a neural network weight matrix ${\bm{W}}^{(l)}$ (on the right) and then is passed through a nonlinear activation function $\sigma(\cdot)$. ${\bm{P}}$ corresponds to the AGGR function and ${\bm{W}}^{(l)}$ and $\sigma(\cdot)$ correspond to the UPDATE function (when it is an MLP) in Equation (1). ${\bm{P}}$ can take different forms for different GNNs, and it is usually a function of ${\bm{A}}$. For example, ${\bm{P}}$ is a degree-normalized version of the adjacency matrix with self-loops for GCN, i.e., ${\bm{P}}=\tilde{\bm{D}}^{-\frac{1}{2}}\tilde{\bm{A}}\tilde{\bm{D}}^{-\frac{1}{2}}$.
Algorithm 1 SAGA (Scatter-ApplyEdge-Gather-ApplyVertex) based GNN Message Passing
  Input: A graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$, node features ${\bm{X}}$
  ${\bm{H}}^{(0)}={\bm{X}}$
  for $l\leftarrow$ 1 to $L$ do
     for (u, v) in ${\mathcal{E}}$ do
        ${\bm{m}}_{u,v}^{(l)}\leftarrow\phi({\bm{h}}_{v}^{(l-1)},{\bm{h}}_{u}^{(l-1)})$
     end for
     for v in ${\mathcal{V}}$ do
        ${\bm{a}}_{v}^{(l)}\leftarrow\textsc{AGGR}(\{{\bm{m}}_{u,v}^{(l)}|u\in{\mathcal{N}}(v)\})$
        ${\bm{h}}_{v}^{(l)}\leftarrow\textsc{UPDATE}({\bm{h}}_{v}^{(l-1)},{\bm{a}}_{v}^{(l)})$
     end for
  end for
The matrix version looks cleaner but it is implemented less often in practice since the full matrix representation of ${\bm{P}}$ usually has space complexity $O(|{\mathcal{V}}|^{2})$, so it is infeasible to fit such ${\bm{P}}$ in memory for large graphs. More frequently, ${\bm{P}}$ is represented as a sparse matrix (e.g., in COO or CSR format) mentioned above, and message passing is implemented using the Scatter-ApplyEdge-Gather-ApplyVertex (SAGA) paradigm (Ma et al., 2019a). For SAGA, messages are first computed and stored on edges, and then each node aggregates messages from its edges and updates the representation. We show the SAGA GNN message passing in Algorithm 1, which has time complexity $O(|{\mathcal{E}}|)$.
2.3. Graph Machine Learning Tasks
GNNs are capable of solving many graph machine-learning tasks. For example, at the node level, it can classify fraudulent nodes. At the edge level, it can predict edges for recommendations. At the graph level, it can classify properties for molecule graphs. For all three kinds of tasks, the standard approaches share a common step of getting node representations ${\bm{H}}$. Then ${\bm{h}}_{v}$ can be used for solving node-level tasks, or pairs $({\bm{h}}_{u},{\bm{h}}_{v})$ can be used for solving edge-level tasks, and ${\bm{H}}$ may go through a pooling layer over all of its rows to become a single representation for solving graph-level tasks. The last prediction step is usually running through a few fully-connected network layers and is not time-consuming (Kipf and Welling, 2016). Therefore, for GNN acceleration, we focus on how ${\bm{H}}$ can be efficiently learned on large graphs, and we use multi-class node classification as a concrete example in this survey.
Node classification is one of the most widely considered tasks on graphs, which classifies each node $v\in{\mathcal{V}}$ to one of $C$ classes. It is usually semi-supervised meaning that the entire graph with all node features and a few node labels are observed. The goal is to leverage the observed labels and the graph information to classify other nodes without labels. When inferring the class of node $v$, its final layer hidden representation ${\bm{h}}_{v}^{(L)}$ is first generated via message passing, and then ${\bm{h}}_{v}^{(L)}$ is plugged into a prediction head $f:{\mathcal{R}}^{F}\rightarrow{\mathcal{R}}^{C}$ to get the final prediction $\hat{\bm{y}}_{v}=f({\bm{h}}_{v}^{(L)})$. $f(\cdot)$ is usually an MLP learned jointly with the GNN using the cross-entropy loss. The loss will be computed for all nodes with observed labels during training, and for a single node $v$, it is:
(3) ${\mathcal{L}}({\bm{y}}_{v},\hat{\bm{y}}_{v})=\sum_{c=1}^{C}{\bm{y}}_{v,c}\log\hat{\bm{y}}_{v,c}.$
3. GNN Acceleration: Training Algorithms
Given a $\textsc{GNN}_{\theta}$ with randomly initialized parameters $\theta$ and a graph ${\mathcal{G}}$ with node features ${\bm{X}}$ and labels ${\bm{Y}}$, the goal of GNN training is to find an optimal set of parameters $\theta^{*}$ that minimizes the loss, i.e., $\theta^{*}=\operatorname*{arg\,min}_{\theta}{\mathcal{L}}(\textsc{GNN}_{\theta}({\mathcal{G}},{\bm{X}}),{\bm{Y}})$. The major latency for GNN training algorithms comes from aggregating messages from (multi-hop) neighbor nodes in the receptive field, and the computation graphs can become huge for deep GNNs on large graphs. The general idea of GNN training acceleration is thus to reduce the computation graphs. Also, training with acceleration is desired to produce a GNN with similar performance to the GNN trained without acceleration.
In this section, we discuss two categories of training acceleration techniques: graph modification and sampling. Both methods aim to reduce the computation graph to accelerate training. The difference is whether there is a modified graph ${\mathcal{G}}^{\prime}$ as an intermediate output. For graph modification, the procedure can be separated into two steps. The first step outputs a graph ${\mathcal{G}}^{\prime}$ that is smaller than ${\mathcal{G}}$ and can be fastly trained on. The second stage is regular GNN training with ${\mathcal{G}}^{\prime}$ as the input. For sampling, a subset of nodes/edges is selected to construct smaller computation graphs in each training iteration. Generally speaking, sampling also modifies ${\mathcal{G}}$, but the modification is done dynamically and implicitly, so there is no intermediate ${\mathcal{G}}^{\prime}$ output. Since the computation graph can differ from iteration to iteration, all the nodes and edges have the chance to be covered by sampling. In contrast, for graph modification, not all the nodes and edges will appear in ${\mathcal{G}}^{\prime}$, but new nodes and edges may be created.
3.1. Graph Modification
Graph modification accelerates GNN training in two steps. The first step takes the graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$, node features ${\bm{X}}$, and labels ${\bm{Y}}$ as input and outputs modified ${\mathcal{G}}^{\prime}=({\mathcal{V}}^{\prime},{\mathcal{E}}^{\prime})$, ${\bm{X}}^{\prime}$, and ${\bm{Y}}^{\prime}$ for training. Training on ${\mathcal{G}}^{\prime}$ will be faster, if ${\mathcal{G}}^{\prime}$ is smaller than ${\mathcal{G}}$ in the sense that $|{\mathcal{V}}^{\prime}|<|{\mathcal{V}}|$ and/or $|{\mathcal{E}}^{\prime}|<|{\mathcal{E}}|$. The second step trains a GNN on ${\mathcal{G}}^{\prime}$ such that the GNN has a similar performance on ${\mathcal{G}}$ as if it is trained on ${\mathcal{G}}$. Mathematically, regular training outputs $\theta^{*}=\operatorname*{arg\,min}_{\theta}{\mathcal{L}}(\textsc{GNN}_{\theta}({\mathcal{G}},{\bm{X}}),{\bm{Y}})$, and training with graph modification outputs $\theta^{\prime*}=\operatorname*{arg\,min}_{\theta^{\prime}}{\mathcal{L}}(\textsc{GNN}_{\theta^{\prime}}({\mathcal{G}}^{\prime},{\bm{X}}^{\prime}),{\bm{Y}}^{\prime})$ such that $\textsc{GNN}_{\theta^{\prime*}}$ has similar performance as $\textsc{GNN}_{\theta^{*}}$ on ${\mathcal{G}}$. In the following, we go over several types of graph modification methods including graph coarsening, graph sparsification, and graph condensation. Each of these methods modifies the graph ${\mathcal{G}}$ differently to create its own ${\mathcal{G}}^{\prime}$, but all ${\mathcal{G}}^{\prime}$ are smaller graphs that accelerate GNN training. An illustration of these graph modification methods is shown in Figure 2.
Figure 2. Illustration of graph modification methods: Graph Coarsening methods perform graph clustering and merge clusters of nodes into a super-node. Graph Sparsification methods remove less important edges. Graph Condensation methods generate a new condensed graph using a randomly initialized generative model. For the modified graph (the rightmost column), black nodes/edges are from the original graph, and colored nodes/edges are newly created.
3.1.1. Graph Coarsening
Graph coarsening is a technique aiming at reducing the graph size while preserving its overall structure. Given a graph ${\mathcal{G}}$, graph coarsening merges nodes in the same local structure of ${\mathcal{G}}$ into a single “super-node” and merges edges connecting super-nodes into a “super-edge”. The coarse graph ${\mathcal{G}}^{\prime}$ thus has much fewer nodes and edges compared to ${\mathcal{G}}$. The key step of coarsening is to cluster nodes in ${\mathcal{G}}$ into $K$ clusters for merging. Nodes in the same cluster will be merged into a super-node, and then a super-edge $(k,o)$ connecting super-node $k$ and super-node $o$ will be constructed if there are edges connecting nodes in cluster $k$ and nodes in cluster $o$. An illustration of the clustering and merging of graph coarsening is shown in Figure 2 top row.
Graph coarsening, with its central step being graph clustering, has been studied long before GNNs were proposed (Kernighan and Lin, 1970; Karypis and Kumar, 1998; Von Luxburg, 2007), and new coarsening algorithms have been proposed until very recently (Loukas and Vandergheynst, 2018; Loukas, 2019; Bravo Hermsdorff and Gunderson, 2019). The key idea of these algorithms is to coarsen the graph while preserving some graph properties, which are often related to the graph spectrum, e.g., restricted spectral approximation (Loukas and Vandergheynst, 2018) and inverse Laplacian (Bravo Hermsdorff and Gunderson, 2019). Below, we discuss the graph coarsening algorithms related to GNNs.
Huang et al. (Huang et al., 2021b) discuss a general framework using graph coarsening to accelerate GNN training. In particular, a graph ${\mathcal{G}}$ is partitioned into K clusters with sizes $N_{1},...,N_{K}$. A matrix ${\bm{C}}\in{\mathcal{R}}^{N\times K}$ is used to represent the partition, where ${\bm{C}}_{i,j}=1/\sqrt{N_{j}}$ if and only if node $i$ is assigned to cluster $j$, and ${\bm{C}}_{i,j}=0$ otherwise, so that ${\bm{C}}^{\top}{\bm{C}}={\bm{I}}$. The super-nodes and super-edges are constructed according to ${\bm{C}}$. Moreover, the super-node features ${\bm{X}}^{\prime}$ are set to be the weighted average of node features within each cluster, i.e., ${\bm{X}}^{\prime}={\bm{C}}^{\top}{\bm{X}}$. The super-nodes labels ${\bm{Y}}^{\prime}$ are constructed by taking the dominant label of nodes in each super-node, i.e., ${\bm{y}}^{\prime}_{k}=\textsc{ONEHOT}(\operatorname*{arg\,max}(({\bm{C}}^{\top}{\bm{Y}})_{k}))$. (Huang et al., 2021b) considers different partition algorithms for coarsening the graph: spectral clustering (Von Luxburg, 2007) and several partition algorithms from (Loukas, 2019) with restricted spectral approximation. Empirically, the Variation Neighborhoods method from (Loukas, 2019) works better than other partition algorithms.
There are some other related works on the topic of graph coarsening and GNNs, but their goals are different from our focus on graph coarsening for GNN training acceleration. GOREN (Cai et al., 2021) uses GNNs as a tool to help do better graph coarsening, where the goal is to output a coarse graph with better Rayleigh loss and Eigenerror. GraphZoom (Deng et al., 2020) uses graph coarsening to improve the accuracy and scalability of graph embedding algorithms like DeepWalK (Perozzi et al., 2014) and Node2Vec (Grover and Leskovec, 2016).
3.1.2. Graph Sparsification
Graph sparsification by removing redundant edges is a natural idea of graph modification to reduce the computation graph and accelerate GNN training. An illustration of graph sparsification is shown in Figure 2 middle row. Like graph coarsening, graph sparsification has been studied and widely used before GNNs became popular (Benczúr and Karger, 1996; Spielman and Srivastava, 2011; Spielman and Teng, 2014; Serrano et al., 2009). Graph sparsification methods usually try to find a sparsified ${\mathcal{G}}^{\prime}$ that can well approximate the properties of the original ${\mathcal{G}}$. That is keeping all the nodes in ${\mathcal{G}}$ and removing less-important edges without disconnecting ${\mathcal{G}}$. Example properties preserved by existing graph sparsification methods include the total weight of cuts (Benczúr and Karger, 1996), spectral properties (Spielman and Srivastava, 2011; Spielman and Teng, 2014), and hierarchical structures (Serrano et al., 2009).
With the goal being GNN training acceleration, the above graph sparsification algorithms can be applied as a pre-processing step before GNN training starts. Since the sparsification algorithms are usually highly scalable and only need to be run once, the time it takes is negligible compared to GNN training. A sparsified graph can reduce the full-batch time complexity of training an $L$-layer GNN from $O(L|{\mathcal{E}}|)$ to $O(L|{\mathcal{E}}^{\prime}|)$, where $|{\mathcal{E}}^{\prime}|$ could be much smaller than $|{\mathcal{E}}|$.
There are some related works performing GNN-based graph sparsification to improve GNN accuracy (Li et al., 2020) or robustness (Zheng et al., 2020b) but not for GNN acceleration explicitly. Their difference from the traditional methods is the sparsification criteria consider both ${\mathcal{G}}$ and the GNN instead of the properties of ${\mathcal{G}}$ only. These criteria are usually formed as optimization problems and solved while training the GNN. For example, (Li et al., 2020) finds a sparsified graph ${\mathcal{G}}^{\prime}$ via the alternating direction method of multipliers. (Zheng et al., 2020b) learns a sparsification strategy by generating sparse k-neighbor subgraphs. It is not guaranteed that these sparsification methods will always accelerate single GNN training. Although the sparse graph reduces ${\mathcal{E}}$ to accelerate message passing and is likely to make the model training converge faster, solving the extra optimization problem takes time and can slow down the training. These methods may be used to accelerate the training of the GNNs after the first one if multiple GNNs need to be trained on the same graph, as for the neural architecture search task.
A closely related concept that may be confused with sparsification is pruning. Some papers use these two terms interchangeably, but conventionally pruning refers to a modification of the neural network model in the deep learning context. Thus, model pruning and graph pruning are totally different concepts with pruning more often referring to the former. We will discuss model pruning in more detail in Section 4.1 for GNN inference acceleration. For graph pruning, generally speaking, it is broader than graph sparsification and includes pruning graph nodes as well. However, since GNN is trained to learn node representations, it is rare to prune nodes in ${\mathcal{G}}$ completely without preserving their information. What is more often done is dynamically exclude different nodes in each training iteration, but each node has a probability to be included in the entire training process. This is the sampling procedure that we will discuss in Section 3.2.
3.1.3. Graph Condensation
Graph condensation generates a completely new graph that preserves the training dynamics of the original graph (Jin et al., 2021, 2022). In other words, the result of training two GNNs with identical architecture on the original graph ${\mathcal{G}}$ and the condensed graph ${\mathcal{G}}^{\prime}$ (the generated new graph) respectively should match. Meanwhile, ${\mathcal{G}}^{\prime}$ can be much smaller than ${\mathcal{G}}$ and thus accelerating GNN training. An illustration of graph condensation is shown in Figure 2 bottom row.
GCond (Jin et al., 2021) matches the training gradients of two GNNs to generate a condensed graph. In particular, a randomly initialized generative model first generates the condensed graph ${\mathcal{G}}^{\prime}=({\mathcal{V}}^{\prime},{\mathcal{E}}^{\prime},{\bm{X}}^{\prime})$ and labels ${\bm{Y}}^{\prime}$, given a condensation ratio $r$, and two GNNs $GNN_{\theta}$ and $GNN_{\theta^{\prime}}$ are trained in parallel using stochastic gradient descent (SGD) on ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ respectively. At each training step, SGD generates gradients $\nabla_{\theta}{\mathcal{L}}(GNN_{\theta}({\mathcal{G}}),{\bm{Y}})$ and $\nabla_{\theta^{\prime}}{\mathcal{L}}(GNN_{\theta^{\prime}}({\mathcal{G}}^{\prime}),{\bm{Y}}^{\prime})$ for updating $\theta$ and $\theta^{\prime}$ respectively. Then gradient matching is performed by tuning the generative model to minimize the distance between $\nabla_{\theta}{\mathcal{L}}(GNN_{\theta}({\mathcal{G}}),{\bm{Y}})$ and $\nabla_{\theta^{\prime}}{\mathcal{L}}(GNN_{\theta^{\prime}}({\mathcal{G}}^{\prime}),{\bm{Y}}^{\prime})$, so that ${\mathcal{G}}^{\prime}$ and ${\bm{Y}}^{\prime}$ better preserve the training dynamics can be generated in the next iteration. For the generative model, since ${\mathcal{E}}^{\prime}$ and ${\bm{X}}^{\prime}$ depend on each other in general graphs, GCond chooses to generate ${\bm{X}}^{\prime}$ first and then construct ${\mathcal{E}}^{\prime}$ from ${\bm{X}}^{\prime}$. Specifically, GCond initializes $r|{\mathcal{V}}|$ node features ${\bm{X}}^{\prime}$ as free learnable parameters, which will be directly updated by the gradient-matching loss. To generate condensed edges ${\mathcal{E}}^{\prime}$, GCond uses an MLP as a feature-based structure prediction model, which takes input ${\bm{x}}^{\prime}_{i}$ and ${\bm{x}}^{\prime}_{j}$ and predicts edge $i$-$j$. Finally, for condensed labels ${\bm{Y}}^{\prime}$, GCond initializes a ${\bm{y}}^{\prime}$ for each ${\bm{x}}^{\prime}_{i}$ and fixes ${\bm{y}}^{\prime}$ by randomly selecting a class following the label distribution in the original graph ${\mathcal{G}}$ without learning. Since all ${\bm{X}}^{\prime}$ are randomly initialized, it doesn’t matter what is the value of ${\bm{y}}^{\prime}$ assigned to each ${\bm{x}}^{\prime}_{i}$, and only the distribution of ${\bm{Y}}^{\prime}$ matters. GCond sets the distribution ${\bm{Y}}^{\prime}$ to follow the original ${\bm{Y}}$ so that the class ratios in ${\mathcal{G}}$ are kept in ${\mathcal{G}}^{\prime}$.
Note that like the GNN-based sparsification methods, GCond can only accelerate training for the second and subsequent GNNs when multiple GNNs need to be trained on the same graph because the gradient matching step requires training a GNN on the original graph ${\mathcal{G}}$. However, GCond is shown to achieve an astonishing less than 1% condensation ratio on several graph benchmarks while maintaining over 95% accuracy as the original GNN, and the same condensed ${\mathcal{G}}^{\prime}$ can be transferred to train GNNs with different architectures to achieve good performance. It is thus a great fit for tasks like neural architecture search. A neural architecture search experiment is performed in GCond, and hundreds of GNNs were quickly trained to achieve competitive performance after a good ${\mathcal{G}}^{\prime}$ is generated. However, to accelerate the training of a single GNN, GCond should either be combined with other acceleration methods or better condensation strategies that can match $GNN_{\theta}$ and $GNN_{\theta^{\prime}}$ more efficiently without training to convergence is needed.
3.2. Sampling
Sampling is a prevalent technique for accelerating GNN training, especially under computation resource limitation (Liu et al., 2021b; Serafini and Guan, 2021). The basic idea is to dynamically select a subset of nodes and edges during each training iteration and build the computation graph only with the selected nodes and edges. Sampling can greatly reduce the computation graph size and accelerate model training. It is found that the model accuracy can also be maintained by sampling since the full neighborhood is not always required to predict a target node. The shared goal of different sampling algorithms is to reduce the time and memory consumption as much as possible while maintaining model accuracy. Note that sampling is different from the graph modification discussed above. Sampling is done dynamically and implicitly and there is no intermediate modified graph output.
In the following, we introduce different sampling algorithms in a unified framework. Throughout the section, we use $\textsc{SAMPLE}(\cdot)$ to denote performing the sampling operation and use $\overline{\cdot}$ on our notations to denote the sampled version. Given each target node $v$, instead of aggregating from all its full neighborhood $\mathcal{N}(v)$ for conducting message passing, sampling selects a subset of neighboring nodes, denoted $\overline{\mathcal{N}}(v)$, which provides the most important information for predicting $v$. A sampled version of GNN message passing in Equation (1) and Equation (2) can approximate the node representations as:
(4) ${\bm{h}}_{v}^{(l)}$ $\approx\textsc{UPDATE}({\bm{h}}_{v}^{(l-1)},\textsc{AGGR}(\{\phi({\bm{h}}_{v}^{(l-1)},{\bm{h}}_{u}^{(l-1)})|u\in\overline{{\mathcal{N}}}(v)^{(l)}\}))$
(5) ${\bm{H}}^{(l)}$ $\approx\sigma(\overline{{\bm{P}}}^{(l)}{\bm{H}}^{(l-1)}{\bm{W}}^{(l)})$
where $\overline{{\mathcal{N}}}(v)^{(l)}$ and $\overline{{\bm{P}}}^{(l)}$ indicate the layer-$l$-sampled neighborhood and its corresponding filter matrix respectively.
The central idea of sampling is to find informative neighborhoods that can construct small but effective computation graphs. Existing sampling algorithms can be categorized into three types according to their input to the $\textsc{SAMPLE}(\cdot)$ operation: 1) node-wise, which takes a single node neighborhood as input and sample within it; 2) layer-wise, which takes all nodes in the same computation graph layer and sample all their neighborhoods jointly; 3) subgraph-wise, which takes a whole graph as input and sample a subgraph as neighborhoods for nodes in the graph. Figure 3 illustrates the workflow of these three types of sampling algorithms. In the following, we discuss the details of sampling algorithms in each category and summarize their pros and cons.
Moreover, we compare the performance of the most popular sampling methods in terms of their run time and accuracy. The results are shown in Table 2. The results are on three commonly used benchmark datasets: Pubmed (Sen et al., 2008b), PPI (Zitnik and Leskovec, 2017), and Reddit (Hamilton et al., 2017b). Pubmed (Sen et al., 2008b) is a citation graph consisting of 19,717 scientific publications pertaining to diabetes from the PubMed database as the graph nodes and 44,338 citations as graph links. Each publication is described with a 500-dimensional TF/IDF weighted word vector feature (Manning, 2009). The task is to classify nodes into one of the three classes each representing one type of diabetes. PPI (Zitnik and Leskovec, 2017) is a graph describing protein-protein interaction consisting of 14,755 nodes and 225,270 edges, each node is described with a 50-dimensional feature vector based on the positional gene set, motif gene sets, and immunological signatures. The task is to classify nodes into one of the 121 classes based on their gene ontology set. Reddit (Hamilton et al., 2017b) is a graph constructed from Reddit posts. Each of the 232,965 nodes corresponds to one Reddit post, and two nodes are connected if the same user comments on both posts, resulting in 11,606,919 links in the graph. The node features are 602-dimensional vectors concatenating the average embedding of the post title and that of all the post’s comments, the post’s score, and the number of comments made on the post, where the average embedding is obtained by averaging the 300-dimensional GloVe CommonCrawl word vectors (Pennington et al., 2014). The task is to classify posts (nodes) into one of the 41 Reddit communities. For all the datasets and all methods, we report their relative and absolute accuracy and total running time. The layer-wise sampling methods work better on Reddit, and the subgraph-wise methods work better on PPI.
Figure 3. Illustration of different Graph Sampling Methods. Node-wise Sampling methods sample per node for each computation layer, which might lead to redundant nodes (e.g., node E is sampled twice), and missing edges (e.g., edge between node C and node D is missing); Layer-wise Sampling methods sample per layer based on nodes in the previous layer. Subgraph-wise Sampling methods sample a list of nodes and their induced subgraph, and then conduct message passing via all the edges within the sampled subgraph.
3.2.1. Node-wise Sampling
Node-wise sampling methods are applied to each target node $v$ at each layer $l$. The node-wise sampled neighborhood is obtained by: $\overline{{\mathcal{N}}}(v)^{(l)}\leftarrow\textsc{SAMPLE}^{(l)}\big{(}\mathcal{N}(v)\big{)}$. Given $s_{node}^{(l)}$ as the number of neighboring nodes we want to sample for each node at layer $l$, the sampling operation $\textsc{SAMPLE}^{(l)}(\mathcal{N}(v))$ is performed $s_{node}^{(l)}$ times. We use the distribution $u\sim p^{(l)}_{v}(u)$ to denote the probability node $u$ is sampled as part of the computation graph for node $v$. An example of the sampling procedure is illustrated in Figure 3(a).
GraphSAGE (Hamilton et al., 2017a) is the pioneering work to reduce node receptive field with node-wise sampling. For each layer $l$, GraphSAGE uniformly samples a fixed number ($s_{node}^{(l)}$) of neighboring nodes for each node. In other words, each neighbor node $u$ in ${\mathcal{N}}(v)$ has an equal probability $1/d(v)$ to be sampled, i.e. $p^{(l)}_{v}(u)=1/d(v)$. The same sampling procedure is repeated multiple times to construct a computation graph layer that is smaller than the full neighbor.
The GraphSAGE sampling algorithm can also be formalized in matrix form, with ${\bm{P}}^{(l)}{\bm{H}}^{(l-1)}$ denoting the full-neighbor aggregation and $\overline{{\bm{P}}}^{(l)}{\bm{H}}^{(l-1)}$ denoting the aggregation with sampling. The sampled filter matrix $\overline{{\bm{P}}}^{(l)}$ can be computed with a mask matrix $\mathcal{M}\in\mathbb{R}^{N\times N}$ as $\overline{{\bm{P}}}^{(l)}=\mathcal{M}\circ{\bm{P}}^{(l)}$. Here $\circ$ means item-wise multiplication, and the mask matrix $\mathcal{M}$ is defined as:
(6) $\mathcal{M}_{u,v}=\begin{cases}\frac{|\mathcal{N}(v)|}{s_{node}^{(l)}}&u\in\overline{{\mathcal{N}}}(v)^{(l)}\ \ \text{, where}\ \ \overline{{\mathcal{N}}}(v)^{(l)}\leftarrow\texttt{SAMPLE}^{(l)}\big{(}\mathcal{N}(v)\big{)}\\ 0&\text{otherwise}\end{cases}$
The sampled estimation $\overline{{\bm{P}}}^{(l)}{\bm{H}}^{(l-1)}$ is an unbiased estimation of ground-truth ${\bm{P}}^{(l)}{\bm{H}}^{(l-1)}$, i.e., the expectation $\mathbb{E}\big{[}\overline{{\bm{P}}}^{(l)}{\bm{H}}^{(l-1)}\big{]}={\bm{P}}^{(l)}{\bm{H}}^{(l-1)}$, as sampling is uniform over each node.
VR-GCN (Chen et al., 2017) proposes to use historical embeddings in GraphSAGE workflow to reduce embedding variance and accelerate training convergence. VR-GCN applies the same sampling operation with the same uniform probability distribution as in GraphSAGE. Under the assumption that the model weight ${\bm{W}}^{(l)}$ would not change significantly between two training iterations, VR-GCN stores the historical embeddings ${\bm{H}}^{(l,\ history)}$ from the previous epoch. With that, ${\bm{P}}^{(l)}{\bm{H}}^{(l-1)}$ is be estimated by $\overline{{\bm{P}}}^{(l)}{\bm{H}}^{(l-1)}+{\bm{P}}^{(l)}{\bm{H}}^{(l-1,\ history)}$. Such historical embedding usage causes some extra memory, but it is usually acceptable. The benefit is that VR-GCN can enjoy the advantage of a smaller training variance that accelerates training convergence. In practice, (Chen et al., 2017) found that VR-GCN can get a comparable performance of GraphSAGE and other existing methods with a much smaller sample size.
Except for GraphSAGE and VR-GCN, there are many other works exploring different aspects of node-wise sampling. Ying et al. (2018a) propose PinSAGE Sampler, which is a node-wise sampling with random walk frequencies-based sampling distribution. It allows GNNs to be applicable to industry-level large-scale recommender systems. MVS-GNN (Cong et al., 2020) extends VR-GCN by explicitly reducing both the embedding approximation variance and stochastic gradient variance. GCN-BS (Liu et al., 2020) formulates the neighbor sampling as a multi-armed bandit problem, and proposes to estimate a learnable sampler that can be updated towards minimal sampling variance. BNS (Yao and Li, 2021) first conducts random node-wise sampling, then uniformly selects a small portion of sampled neighbors to stochastically blocks their ongoing expansion.
Although node-wise sampling methods can alleviate the scalability challenge of large-scale GNN training by reducing the receptive field and can guarantee that each node at each layer can have a reasonable number of connected neighbors to update its hidden representation, they still have some limitations:
  • The exponential neighbor growth is not avoided as the GNN goes deeper, even though the neighborhood size is reduced. As shown in Figure 3(a), the computation graph keeps growing by a factor of size $s_{node}^{L}$ in each layer.
  • A node can be repeatedly sampled since the sampling procedures for each node in the same layer is independent. In Figure 3(a), node E is sampled by both node B and node C, and the message passing computation is repeated twice in the computation graph (marked by the blue dotted box).
  • Some edges between sampled nodes across layers are missing, which leads to computation resource waste. In Figure 3(a), node C and node D are connected in the original graph. However, since node D is sampled by node B but not node C, the edge (B, C) is missing, which could have been used to pass messages from B to C given these two are included in the computation graph anyway.
Table 2. Performance comparison of different sampling methods. Absolute performance is adopted from (Liu et al., 2021b), and we add relative performance with GraphSAGE set as the benchmark (1$\times$), as it is the simplest and the most widely used sampling method. All results assume a 2-layer backbone GCN model with the sampling size and the optimizer the same as the original papers.
Pubmed PPI Reddit
Accuracy (%) Total Time (s) Accuracy (%) Total Time (s) Accuracy (%) Total Time (s)
No Sample GCN 78.6 (0.97$\times$) 24.8 (0.65$\times$) N/A N/A N/A N/A
Node-wise GraphSAGE 81.4 (1$\times$) 37.9 (1$\times$) 61.6 (1$\times$) 47.5 (1$\times$) 95.0 (1$\times$) 326.7 (1$\times$)
VR-GCN 81.4 (1$\times$) 5.9 (0.16$\times$) 97.5 (1.58$\times$) 83.5 (1.76$\times$) 96.3 (1.01$\times$) 390.6 (1.20$\times$)
Layer-wise FastGCN 86.3 (1.06$\times$) 13.7 (0.36$\times$) N/A N/A 92.8 (0.98$\times$) 294.5 (0.90$\times$)
AS-GCN 89.5 (1.10$\times$) 39.1 (1.03$\times$) N/A N/A 96.5 (1.02$\times$) 1633.7 (5.00$\times$)
LADIES 75.7 (0.93$\times$) 6.4 (0.17$\times$) N/A N/A N/A N/A
Subgraph-wise Cluster-GCN N/A N/A 95.8 (1.56$\times$) 202.3 (4.26$\times$) 95.8 (1.01$\times$) 981.8 (3.00$\times$)
GraphSAINT N/A N/A 98.7 (1.60$\times$) 550.1 (11.58$\times$) 96.7 (1.02$\times$) 124.0 (0.38$\times$)
3.2.2. Layer-wise Sampling
Layer-wise sampling methods are applied to all nodes in the same computation layer instead of a single node. When a set of nodes are selected as the prediction target during training, a computation graph is constructed with all these nodes belonging to the first computation layer. Their direct neighbors belong to the next computation layer, and their 2nd-hop neighbors belong to the computation layer after, and so on. Let ${\mathcal{V}}^{(l)}$ denote all the nodes at the $l-$th layer, layer-wise sampling methods choose a subset of nodes from ${\mathcal{V}}$ to be in the next computation layer. In this way, for each node $v\in\mathcal{V}^{(l)}$, its layer-sampled neighborhood is the intersection of the whole layer-sampled node set and its original neighborhood: $\overline{{\mathcal{N}}}(v)^{(l)}\leftarrow\texttt{SAMPLE}({\mathcal{V}})\cap\mathcal{N}(v)$. Let $s_{layer}^{(l)}$ be the number of neighbor nodes to be sampled for $l$, each Layer-wise sampling operation $\textsc{SAMPLE}^{(l)}({\mathcal{V}})$ is performed $s_{layer}^{(l)}$ times. Finally, the sampled nodes at the current computation layer $l$ would be used to construct the neighborhood of the later computation layer $l+1$ for message passing. An example of the sampling procedure is illustrated in Figure 3(b).
FastGCN (Chen et al., 2018a) proposes the first layer-wise importance sampling scheme to address the scalability issue in node-wise sampling. For each computation layer, FastGCN samples nodes from the full node set ${\mathcal{V}}$ with the node-degree-based probability distribution: $p^{(l)}_{\mathcal{V}}(u)=\|{\bm{P}}_{:,u}\|^{2}_{2}/\sum_{u^{\prime}\in\mathcal{V}}\|{\bm{P}}_{:,u^{\prime}}\|^{2}_{2}$. Note that, in FastGCN, the same sampling distribution is used for different layers, i.e. $p^{(l)}_{\mathcal{V}}(u)=p_{\mathcal{V}}(u)$ for all $l$. Next, the approximation of ${\bm{P}}^{(l)}{\bm{H}}^{(l-1)}$ with $\overline{{\bm{P}}}^{(l)}{\bm{H}}^{(l-1)}$ can be computed by constructing $\overline{{\bm{P}}}^{(l)}$. The $(u,v)$ entry of $\overline{{\bm{P}}}^{(l)}$ is given by
(7) $\overline{{\bm{P}}}^{(l)}_{v,u}=\begin{cases}\frac{{\bm{P}}_{v,u}^{(l)}}{s_{layer}^{(l)}p_{\mathcal{V}}(u)}&\text{if }v\text{ is sampled in layer ($l$) and }u\in\mathcal{N}(v)\\ 0,&\text{otherwise}\end{cases}$
Despite its efficacy and efficiency, FastGCN suffers the connectivity challenge since the sampling for each layer is conducted independently. In other words, the nodes sampled in two consecutive layers may not be connected, some nodes sampled in the previous layer may not be able to pass their information to the next layer, and some nodes sampled from the later layer may not be able to get any information from its neighborhood. This would lead to a very sparse $\overline{{\bm{P}}}^{(l)}$ matrix and may make the approximated embeddings have a large variance.
AS-GCN (Huang et al., 2018) proposes an adaptive layer-wise sampling method to accelerate the training of GCNs on large-scale graphs. The proposed method samples the lower layer conditioned on the top one, where the sampled neighborhoods are shared by different parent nodes and over-expansion is avoided. AS-GCN explicitly calculate the embedding variance, and add a regularization to reduce the variance. In this way, the model could learns to give higher weights to those more connected nodes, which has been demonstrated to achieve higher accuracy than FastGCN. Additionally, AS-GCN preserves second-order proximity by formulating a skip connection across two layers.
LADIES (Zou et al., 2019) adopts similar intuition as AS-GCN to only sample within the the immediate neighborhood of the later layer. This constraint can guarantee that each sampled node in the previous layer will have at least one successor node in the later layer. In other words, it can make sure all the sampled nodes can deliver their information to at least one of their neighbor nodes in the next layer. To apply this constraint, a row selection matrix ${\bm{Q}}^{(l)}\in{\mathcal{R}}^{s_{layer}^{(l)}\times|{\mathcal{V}}^{(l)}|}$ is constructed as
(10) ${\bm{Q}}^{(l)}_{v,u}=\left\{\begin{array}[]{ll}1&\text{if }v\text{ is sampled in layer $l$ and }u\in\mathcal{N}(v)\\ 0,&\text{otherwise},\end{array}\right.$
Then, node $u$ in layer $(l-1)$ is sampled with probability: $p^{(l-1)}_{\mathcal{V}}(u)=\frac{\|{\bm{Q}}^{(l)}{\bm{P}}_{:,u}\|^{2}_{2}}{\|{\bm{Q}}^{(l)}{\bm{P}}\|^{2}_{F}}$. A $\overline{{\bm{P}}}^{(l)}$ can be constructed in the same way as in Equation (7).
To maintain the scale of embeddings in the forward process, LADIES also normalizes $\overline{{\bm{P}}}^{(l)}$ to compute the $\overline{{\bm{P}}}^{(l)}{\bm{H}}^{(l-1)}$. It computes $\overline{{\bm{P}}}^{(l)}\leftarrow{\bm{D}}_{\overline{{\bm{P}}}^{(l)}}^{-1}\overline{{\bm{P}}}^{(l)}$, where ${\bm{D}}_{\overline{{\bm{P}}}^{(l)}}$ is a diagonal matrix with its $(i,i)$ entry equals $\sum_{j}\overline{{\bm{P}}}^{(l)}_{ij}$.
Layer-wise sampling methods solve the exponential neighborhood expansion limitation for node-wise sampling methods and enjoy a linear time and memory complexity with respect to the number of GNN layers. However, the connectivity problem still exists despite the efforts to make the sampled nodes in two consecutive layers connected and guarantee all the sampled nodes can pass out their information. The connectivity problem says that sampled nodes in a later layer can have very limited or no sampled neighbors from the previous layer. Thus, the messages passed between two consecutive layers for some sampled nodes can be insufficient, which degrades the model performance.
3.2.3. Subgraph-wise Sampling
Subgraph-wise sampling methods are applied with the whole graph as input and output a sample subgraph, i.e., $\overline{{\mathcal{G}}}\leftarrow\textsc{SAMPLE}({\mathcal{G}})$. Afterward, GNN training takes the sampled subgraph as input to conduct full-batch message passing. For these methods, the sampling procedure may be conducted before the training starts as a pre-processing step, which makes them more similar to graph modification methods. However, there are still two key differences that make these methods sampling rather than modification. First, to overcome the scalability challenge, the subgraphs are usually much smaller than the original, which only represent local information instead of global information as the graph modification methods discussed in Section 3.1. Second, directly training with such pre-processed small subgraphs often result in low model accuracy. Thus, some stochasticity needs to be added back during training to improve accuracy, which makes the modification in these methods dynamic, e.g., the ClusterGCN method below.
ClusterGCN (Chiang et al., 2019) utilizes graph clustering algorithms (i.e., METIS (Karypis and Kumar, 1998)) to partition the input whole graph $\mathcal{G}$ into clusters, where the nodes within each cluster form a densely connected subgraph and inter-cluster edges are minimized. Afterward, ClusterGCN conducts full-batch GNNs training on each subgraph cluster, which significantly improves the scalability of GNN training. One limitation of this approach is that the clustering result is fixed, so inter-cluster edges will be missing throughout training. ClusterGCN further proposes to randomly choose several subgraphs in each training mini-batch and add back inter-cluster edges between the subgraphs. This operation covers all the inter-cluster edges in the long run and is proven to provide an unbiased estimate of the full-neighborhood embeddings.
GraphSAINT (Zeng et al., 2019) is another widely used subgraph-wise sampling method, which first samples nodes and then constructs subgraphs induced by the sampled nodes. GraphSAINT proposes four different algorithms to sample nodes: 1) node sampler, which randomly samples nodes with probability proportional to node degrees; 2) edge sampler, which samples edges with probability proportional to the inverse of the end node degrees, and then keeps the end nodes. The edge sampler is proven to reduce the estimation variance; 3) random walk sampler, which conducts random walks on a random set of nodes and selects nodes on the walks; 4) multi-dimensional random walk, which is similar to random walk but adds an additional requirement to start random walking from high-degree nodes in the sampled set. Among the four proposed algorithms, the random-walk sampler performs better empirically and thus is more widely used.
Shadow-GNN (Zeng et al., 2021a) samples different subgraphs for each target node (similar to ego-networks). The sampling could be done via random walk or personalized page rank. Specifically, the authors introduce a design principle that separates the depth of a GNN from its receptive field size, which is the range of nodes and edges that influence a node’s representation. By doing so, the depth of a GNN can be increased without increasing its receptive field size exponentially. It also presents a theoretical analysis of expressivity from three different perspectives, and also rich design components (e.g., subgraph extraction functions, architecture extensions) to implement such design principles.
There also exist many other follow-up works that conduct subgraph-wise sampling. RWT (Bai et al., 2021) proposes a ripple walk sampler. It first randomly samples a small set of nodes, then it randomly samples nodes from the neighbors of nodes in the set and adds them to the set. The process is repeated until a prespecified budget is filled. Zeng et al. (2021b) provides the parallelized version of the existing frontier sampling which can sample subgraphs that can approximate the original graph w.r.t. various connectivity measures. In addition, some layer-wise sampling methods could be used to construct subgraphs (Hu et al., 2020a).
The key advantage of subgraph-wise sampling methods is that they do not dependent on the GNN model and the generated embedding. Therefore, they could be done purely in parallel or beforehand the training procedure. However, this advantage could also be regarded as a limitation, as the sampling only considers the graph structure without taking into account model training dynamics. Therefore, it still remains an open question how to incorporate some variance-reduction studies, including VR-GCN (Chen et al., 2017) and AS-GCN (Huang et al., 2018), into the subgraph-wise sampling procedure.
4. GNN Acceleration: Inference Algorithms
Given ${\mathcal{G}}_{train}$ as the training graph, ${\mathcal{G}}_{test}$ as the test graph, and $\textsc{GNN}_{\theta^{*}}$ as the optimal GNN trained on ${\mathcal{G}}_{train}$, GNN inference generates representations for nodes in ${\mathcal{G}}_{test}$ and uses them to make predictions. The goal of GNN inference acceleration is to construct another model $\tilde{\textsc{GNN}}_{\tilde{\theta}}$ such that $\tilde{\textsc{GNN}}_{\tilde{\theta}}$ inference on ${\mathcal{G}}_{test}$ has similar accuracy as $\textsc{GNN}_{\theta^{*}}$ but is much faster. During the construction of $\tilde{\textsc{GNN}}_{\tilde{\theta}}$, access to ${\mathcal{G}}_{test}$ should not be assumed, since fast inference is desired on unseen test graphs for real-world use cases. In contrast, acceleration methods usually assume access to ${\mathcal{G}}_{train}$, since some extra work like retraining is usually needed for a faster $\tilde{\textsc{GNN}}_{\tilde{\theta}}$ to achieve similar accuracy as $\textsc{GNN}_{\theta^{*}}$. In practice, ${\mathcal{G}}_{test}$ and ${\mathcal{G}}_{train}$ can have substantial overlap or even be the same graph (called the transductive setting). For the sake of generality, we assume they are different graphs in our discussion below.
GNN inference acceleration can be quite different from GNN training acceleration. Training acceleration is more graph-centric and inference acceleration is more model-centric. For training acceleration, the GNN architecture is fixed and only the process for finding $\theta^{*}$ is accelerated. Methods focus on modifying the graph or sampling from the graph to reduce the computation graph and thus accelerate training. For inference acceleration, $\tilde{\textsc{GNN}}_{\tilde{\theta}}$ can be substantially different from $\textsc{GNN}_{\theta^{*}}$ in terms of model architecture, parameter numerical precision, and parameter values. Changes are made only on the model. The test graph ${\mathcal{G}}_{test}$ is assumed unseen and thus untouched. Only for cases where ${\mathcal{G}}_{test}$ and ${\mathcal{G}}_{train}$ are overlapped or being the same, some graph-centric training acceleration methods like graph sparsification can also accelerate inference because the sparsification is carried from ${\mathcal{G}}_{train}$ to ${\mathcal{G}}_{test}$. Even in these cases, graph modification methods like coarsening and condensation do not fit because they can affect the node for inference and thus lose the prediction target. Sampling methods are also rarely used because usually only one forward computation is performed in inference rather than many times in training. Therefore, sampling only a random partial neighborhood, in general, can guarantee neither prediction accuracy nor consistency.
Fast inference with $\tilde{\textsc{GNN}}_{\tilde{\theta}}$ does not mean $\tilde{\textsc{GNN}}_{\tilde{\theta}}$ can be constructed fast. Rather, the speed of $\tilde{\textsc{GNN}}_{\tilde{\theta}}$ construction is often sacrificed to achieve inference acceleration. Nevertheless, as long as a $\tilde{\textsc{GNN}}_{\tilde{\theta}}$ can make fast inference is constructed, the goal is achieved. Even slowly constructed $\tilde{\textsc{GNN}}_{\tilde{\theta}}$ will be useful for many practical scenarios. Especially when the construction doesn’t need to be done very frequently, but inference latency is a big concern. In this section, we discuss three kinds of methods for inference acceleration: pruning, quantization, and distillation. All three methods are widely used for accelerating general DNNs. The works we discuss below adapt them on GNNs for graphs. The major challenge they solve is how to preserve the graph structure information and solve the major source of latency caused by message aggregation.
4.1. Pruning
Model pruning or simply pruning is a popular approach for accelerating neural network (NN) inference (Han et al., 2015b). The research of NN pruning started in the 1990s (Reed, 1993), with the idea being removing unimportant parts of an NN while maintaining its accuracy. The most common NN pruning method is the weight magnitude pruning, which prunes the connections between the NN “neurons" with small weights or equivalently reduces the L1 norm of NN weights. Like other inference acceleration methods, pruning provides a trade-off between model speed and accuracy. When more weights are removed from the NN, the model is likely to become faster but less accurate, and vice versa.
Pruning methods have been proposed for accelerating GNN inference (Zhou et al., 2021b; Chen et al., 2021b). Notice that GNN pruning targets the model and is different from graph pruning which targets the graph data. Graph pruning is a concept related to graph sparsification, which we discussed in Section 3.1.2. For GNN model pruning, selecting which GNN weights to prune is the central question and is nontrivial. It is challenging because GNN inference on a graph node depends on its local graph structure and neighboring nodes, which is more complicated than the regular NN inference on data types with independent instances like image classification.
Zhou et al. (Zhou et al., 2021b) propose to accelerate GNN inference with channel pruning, where channels refer to the dimensions of the node representations ${\bm{h}}$, or equivalently the rows of the GNN weight matrix ${\bm{W}}$. Given a GNN with weights ${\bm{W}}$, Zhou et al. formulate the pruning problem as a LASSO regression problem for each GNN layer, and the objective is:
(11) $\operatorname*{arg\,min}_{\hat{\bm{\beta}},\hat{\bm{W}}}\lVert{\bm{H}}^{(l)}-{\bm{P}}{\bm{H}}^{(l-1)}(\hat{\bm{\beta}}\odot\hat{\bm{W}}^{(l)})\rVert_{2}+\lambda\lVert\hat{\bm{\beta}}\rVert_{1}.$
There are two sets of parameters to optimize for this objective: a learnable mask $\hat{\bm{\beta}}\in{\mathcal{R}}^{F}$ that chooses channels to prune from $F$ total channels, and an updated GNN weight matrix $\hat{\bm{W}}$ for better accuracy of the pruned model. $\hat{\bm{\beta}}$ is initialized as all ones and $\hat{\bm{W}}$ is initialized as ${\bm{W}}$. $\odot$ denotes an element-wise product for a column vector on each column of a matrix. ${\bm{P}}$ denotes the filter matrix for message propagation. Zhou et al. optimize these two sets of parameters alternatively. $\hat{\bm{\beta}}$ is optimized first with fixed $\hat{\bm{W}}$ and stochastic gradient descent (SGD). Then $\hat{\bm{W}}$ can be computed in closed-form as a least square solution for fixed $\hat{\bm{\beta}}$. The L1 regularization shrinks entries of $\hat{\bm{\beta}}$ to zeros, which leads to the pruning of corresponding channels. Zhou et al. show that the GNN accuracy only decreases marginally when $3/4$ of the GNN channels are pruned, which achieves 3.27× speedup on large-scale graph inference.
Another recent popular line of pruning research is the Lottery Ticket Hypothesis (LTH) (Frankle and Carbin, 2018), which shows that a pruned sparse NN can be retrained to achieve similar accuracy as the original NN. The significance of LTH compared to the regular pruning methods is in the retraining. Most pruning methods result in one set of NN weights with minor accuracy loss, but they cannot guarantee that the pruned NN can be retrained with the original ML objective and still maintain similar accuracy loss on the same dataset. LTH shows that the same level of accuracy loss can be achieved with iterative magnitude pruning (IMP), i.e., prune the small NN weights iteratively and perform retraining after each pruning.
Chen et al. (Chen et al., 2021b) test LTH for GNNs by doing IMP. They indeed observe speed gain and marginal accuracy loss even for retrained GNNs. You et al. (You et al., 2021) further test the early-bird LTH on GNNs, which means the retraining in each IMP iteration can stop very early before it converges. Notice that pruning accelerates model inference but not always model training. The iterative pruning requires retraining the model multiple times, possibly resulting in a slower training procedure than before. The advantage of early-bird LTH is to save iterative training time for finding the optimal pruned model. Moreover, in (Chen et al., 2021b) and (You et al., 2021), the input graph is pruned together with the GNN model, so parts of these methods are doing graph pruning (sparsification) as well.
4.2. Quantization
Quantization is a widely used technique for accelerating general ML models. The idea is to use lower numerical precision for the model parameters, say replace the 32-bit floating point numbers (FP32) with 8-bit integers (INT8). For DNNs, a significant part of latency comes from matrix multiplications which boil down to multiply–accumulate (MAC) operations. Quantized low-precision DNNs have a smaller number of MAC operations and thus accelerate inference. Besides the MAC operation reduction, quantization also comes with other benefits that further accelerate the model, like reducing memory access time (Hubara et al., 2018; Gupta et al., 2015). Another reason for quantization to become popular is that it is easy to use. The numerical precision can be easily changed for any given model. The only concern is that blindly reducing the numerical precision of a given model at the inference stage may result in a large accuracy loss. Quantization research is thus aiming for inference acceleration while maintaining inference accuracy as much as possible, and many quantization techniques for achieving this goal have been proposed (Krishnamoorthi, 2018; Han et al., 2015a). Below, we first introduce the idea of quantization on general NNs and then get to quantization for GNNs.
[…]
(12) ${\bm{x}}_{q}=\lfloor\frac{{\bm{x}}}{s}+z\rfloor.$
The quantized ${\bm{x}}_{q}$ is in the desired precision. A dequantization is then performed to map ${\bm{x}}_{q}$ to $\hat{\bm{x}}$:
(13) ${\bm{x}}\approx\hat{\bm{x}}=({\bm{x}}_{q}-z)s.$
With the quantize and then dequantize operations, a $q$-bit representation $\hat{\bm{x}}$ can be used to replace model tensor ${\bm{x}}$. $\hat{\bm{x}}$ closely approximates ${\bm{x}}$, but computation using $\hat{\bm{x}}$ will be much faster.
With the same general quantization idea, there are in general two types of quantization algorithms for NNs: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT) (Nagel et al., 2021; Esser et al., 2019). PTQ takes a trained high-precision NN and directly converts it to lower precision. PTQ is easy to use and can be applied to any architecture without accessing the training pipeline or any data. PTQ mainly focuses on two questions to trade off accuracy for speed: 1) Which tensor object should be quantized, e.g., NN weights or NN activations? 2) What is a proper quantization bit $q$ and a proper range $[q_{min},q_{max}]$? The concern of PTQ is that when the precision $q$ is set low, like 4 bits and below, the accuracy can always be low no matter which tensor or range is chosen. QAT algorithms, on the other hand, model the quantization error during training. For example, train a model with low-precision along with the original high-precision model and minimize the error between them. QAT can thus find quantized models better than PTQ in general. The cost of QAT is that the model training time is longer and data is required.
Existing quantization methods are mostly for convolutional neural networks (CNNs). Quantization for GNNs, however, has its own challenges and requires special treatment. We now discuss GNN quantization methods for both PTQ and QAT.
SGQuant (Feng et al., 2020) is one of the earliest work studying GNN quantization. It is a PTQ method and focuses on the two core PTQ questions mentioned above. For the first question about the quantization object, SGQuant focuses on the quantization of node representations learned by each GNN layer, i.e., $\{{\bm{H}}^{(l)}\}_{1\leq l\leq L}$. SGQuant makes the empirical observation that node representations consume much more memory than GNN weights ${\bm{W}}$. This observation and object choice is unique to GNNs but not for other NNs because of node dependency. Unlike other NNs which process one data instance one at a time, for one target node, GNNs store node representations for all the multi-hop neighboring nodes within the receptive field of the target, which could involve a huge number of nodes. For the second question, SGQaunt focuses on selecting the quantization bit $q$ and simply uses the full range of $q$-representable numbers. SGQuant proposes multi-granularity quantization, where different objects are associated with different bits to optimize quantization efficiency. Applying this principle to graph data, SGQuant uses topology-aware and layer-aware quantization to group node representations, where nodes with similar degrees are grouped together, and representations from the same layer are grouped together. This is essentially grouping rows of $\{{\bm{H}}^{(l)}\}_{1\leq l\leq L}$, where each row corresponds to one node and $l$ corresponds to layer depth. Then node representations are assigned with different bits depending on which group they are from. The multi-granularity quantization decides the bits for each group by optimizing quantized model accuracy over a set of precision candidates.
Degree-Quant (Tailor et al., 2020) is a QAT method quantizes both GNN weights ${\bm{W}}$ and intermediate node representations ${\bm{H}}$, which results in INT8 GNNs with comparable accuracy as their FP32 counterparts but up to 4.7X faster. Degree-Quant is like many other QAT methods for general NNs, for which a quantized model is trained via gradients by simulating its numerical errors from the original model. During the backward pass, the straight-through estimator (STE) (Bengio et al., 2013) is used to compute gradients since the quantization operation, e.g., round to integers, can be non-differentiable. The unique challenge for GNN quantization via STE is that the STE performance is degraded for high-degree nodes, which leads to poor weight updates. Degree-Quant verifies the degradation both theoretically and empirically on common GNNs architectures including GCN, GIN and GAT. It then proposes a protective mask to improve weight update accuracy. A protective mask $\bm{\beta}$ is a length-$N$ binary vector for all $N$ nodes in graph ${\mathcal{G}}$. During the quantization-aware GNN training, node representation ${\bm{h}}_{i}$ with $\bm{\beta}_{i}=1$ is kept with full precision, while ${\bm{h}}_{i}$ with $\bm{\beta}_{i}=0$ and all weights ${\bm{W}}$ are quantized. The protective mask $\bm{\beta}$ is sampled as $\bm{\beta}\sim Bernoulli({\bm{p}})$ with ${\bm{p}}$ containing probabilities for each node. Degree-Quant treats the highest probability and lowest probability in ${\bm{p}}$ as two hyperparameters, and probabilities are assigned to nodes by ranking nodes according to their degrees and interpolating these two values according to the ranking. High-degree nodes are assigned higher probabilities, which are encouraged to be kept by the protective mask for more accurate gradient updates. Degree-Quant also proposes to clip the 0.1% top and bottom quantized values to improve model accuracy, because large fluctuations in the variance of message-passing aggregation are observed.
Overall, quantization is one of the easiest GNN acceleration methods to apply but shows significant speed improvement. It can also be combined with other acceleration methods to further boost speed gain. For example, (Zhao et al., 2020) combines quantization with network architecture search (NAS). In (Zhao et al., 2020), a quantization search space are specified and the best quantized GNN is selected via NAS.
4.3. Distillation
Knowledge distillation (KD) or simply distillation is a technique to compress classification models to smaller and thus faster models (Buciluǎ et al., 2006). It was shown to be able to produce much faster models with very little, if not none, accuracy loss for image classification and speech recognition (Hinton et al., 2015). The basic idea of KD is to train two models in a teacher-student fashion. The teacher outputs a probability vector for each input data containing probabilities for the data to be classified into $C$ classes. The probability vectors (or a re-normalized version of them) are in $[0,1]^{C}$ and are called soft labels, in contrast to the ground-truth hard labels in $\{0,1\}^{C}$. Such soft labels will be used to train a simpler student model. Specifically, the standard KD loss on a data instance $i$ is computed as a distance between the teacher model output (soft label) of data $i$, i.e., $\hat{\bm{y}}_{i}^{t}$, and the student model output of data $i$, i.e., $\hat{\bm{y}}_{i}^{s}$. Two outputs are matched by minimizing this loss, and thus knowledge is transferred from the teacher to the student. Since both outputs are probability vectors, a common choice of the distance metric is the Kullback-Leibler (KL) divergence denoted as $D_{KL}$,
(14) $D_{KL}(\hat{\bm{y}}_{i}^{t}||\hat{\bm{y}}_{i}^{s})=\sum_{j}\hat{\bm{y}}_{ij}^{t}\log\frac{\hat{\bm{y}}_{ij}^{t}}{\hat{\bm{y}}_{ij}^{s}}.$
In most cases, the teacher model can be trained alone first, and the student is trained next with the trained teacher fixed. Thus, we can drop the constant $\hat{\bm{y}}_{ij}^{t}\log\hat{\bm{y}}_{ij}^{t}$ term to get the form after $\propto$ for student training. The knowledge distillation loss $\mathcal{L}_{i}^{KD}$ will only be used to update the parameters of the student but not the teacher,
(15) $\mathcal{L}_{i}^{KD}\propto D_{KL}(\hat{\bm{y}}_{i}^{t}||\hat{\bm{y}}_{i}^{s})=-\sum_{j}\hat{\bm{y}}_{ij}^{t}\log\hat{\bm{y}}_{ij}^{s}.$
KD can be naturally applied in a semi-supervised setting, a common setting for graph node classification, by training the teacher with labeled data and generating soft labels for the unlabeled data to train the student (Xie et al., 2020). KD has been wildly used for compressing and accelerating DNNs like CNNs (Hinton et al., 2015), and recently it has been extended to GNNs as well (Yang et al., 2020; Yan et al., 2020b; Zhang et al., 2022a). We discuss these GNN KD methods in detail below. They mostly follow the KD framework discussed above and mainly differ in two perspectives: 1) What are the teacher and the student models? 2) What is the KD objective?
LSP (Yang et al., 2020) is the pioneer work exploring KD for accelerating GNNs. Targeting node classification, the teacher model LSP considers is a GAT and the student model is a GAT with fewer parameters. LSP proposes a new KD objective termed local structure preserving loss. The intuition is that node representations are generated by aggregating messages within the local structure, so the desired distillation function should transfer not only the final aggregation result but also the local structure as part of the knowledge. LSP defines the local structure as the similarity between each node to its neighbors and encourages the local structure of the student to match the teacher. Specifically, the local structure of node $i$ is represented as a vector ${\bm{s}}^{(i)}\in{\mathcal{R}}^{d(v)}$, with ${\bm{s}}^{(i)}_{j}$ representing the local influence of node $j$ to node $i$ defined as
(16) ${\bm{s}}^{(i)}_{j}=\frac{\exp(\mathcal{SIM}({\bm{h}}_{i},{\bm{h}}_{j}))}{\sum_{k:(k,i)\in{\mathcal{E}}}\exp(\mathcal{SIM}({\bm{h}}_{i},{\bm{h}}_{k}))},\text{ where }\mathcal{SIM}({\bm{h}}_{i},{\bm{h}}_{k})=||{\bm{h}}_{i}-{\bm{h}}_{k}||_{2}^{2}.$
Plug-in node representations ${\bm{h}}$ of the teacher or the student gives the teacher local structure ${\bm{s}}^{(i),t}$ and the student local structure ${\bm{s}}^{(i),s}$ respectively. Then, the local structure preserving loss $\mathcal{L}_{i}^{LSP}$ is defined using the KL divergence, similarly as the $\mathcal{L}_{i}^{KD}$ in Equation (17). The difference being $\mathcal{L}_{i}^{LSP}$ computes the KL-divergence of ${\bm{s}}^{(i),s}$ against ${\bm{s}}^{(i),t}$, instead of the teacher probability output $\hat{\bm{y}}_{i}^{t}$ against student output $\hat{\bm{y}}_{i}^{s}$. That is,
(17) $\mathcal{L}_{i}^{LSP}=D_{KL}({\bm{s}}^{(i),s}||{\bm{s}}^{(i),t})=\sum_{j\in{\mathcal{N}}(i)}{\bm{s}}^{(i),s}_{j}\log\frac{{\bm{s}}^{(i),s}_{j}}{{\bm{s}}^{(i),t}_{j}}.$
TinyGNN (Yan et al., 2020b) shares a similar intuition as LSP to preserve local structures, but it chooses to add a special layer termed a Peer-Aware Module (PAM) to the student model instead of defining a new KD objective as in LSP. TinyGNN adopts the standard KL divergence loss $\mathcal{L}_{i}^{KD}$. The teacher model TinyGNN considers is GAT. The student model, however, is a GAT with fewer message-passing layers than the teacher but a special PAM layer inserted before each message-passing layer. The idea of PAM is to encourage nodes in the same message-passing hierarchy to directly interact. For example, when the GNN is only one layer, a message passing step centered at node $v$ will aggregate messages from its neighbors ${\mathcal{N}}(v)$, but nodes in ${\mathcal{N}}(v)$ will not directly interact with each other during this message passing step. In other words, the computation graph is a one-layer tree and has no edges between its leaves. TinyGNN argues that sharing messages between nodes in ${\mathcal{N}}(v)$ is important for capturing the local structure around $v$. This message sharing happens naturally when the GNN is deep like the teacher, but it is lacking if the GNN is only one or two layers like the student in TinyGNN. Thus, the PAM is proposed to encourage this message sharing by adding a self-attention layer between all nodes in ${\mathcal{N}}(v)$. It was shown that a GAT student with PAM has a higher distilled model accuracy than a regular GAT student, though the speed gain is not as much given the extra self-attention computation.
GLNN (Zhang et al., 2022a) explores KD for accelerating GNNs by making the student model as simple as a “graph-less neural network”, i.e., an MLP. GLNN follows the standard KL divergence loss $\mathcal{L}_{i}^{KD}$, but the student model in GLNN is a pure MLP (a zero-layer GNN) with no node dependency. GLNN shows that even with a student as simple as an MLP, KD can help the student MLP to achieve competitive performance to the teacher GNN on many benchmark datasets, even when used for the inductive prediction on unseen new nodes. GLNN can accelerate GNNs to orders of magnitude faster because the node dependency is dropped for the student. It is also shown to work with different teacher GNNs including GCN, GAT, GraphSAGE, etc. However, GLNN is still not a universal solution for all cases. The analysis in (Zhang et al., 2022a) shows that GLNN fails to work when the conditional mutual information between the node features ${\bm{X}}$ and the node labels ${\bm{Y}}$ given the graph structure ${\bm{A}}$ is low. Intuitively, if node labels depend mainly on the graph structure but not the node feature, then MLPs won’t be able to classify the nodes even with KD.
Overall, the methods above apply the general KD idea to GNNs by adapting graph properties different perspectives, including both new objects and new student designs. They all show exciting improvements over the original teacher GNN, but this line of research is still at an early stage. These KD methods were developed under different settings. Thus, we are unaware of the applicability of these distillation techniques to each others setting and whether we can combine them. Since KD in general is always a trade-off between model accuracy and model speed, we don’t know which method or which combination will give the best result balancing these two perspectives. Also, we don’t know under what exact conditions, simple methods like GLNN (Zhang et al., 2022a) will be applicable. These many interesting open questions make KD a promising direction to explore.
Finally, there are some other related works applying KD techniques on GNNs, for which the goal is less accuracy loss but not necessarily inference acceleration (Yang et al., 2021b; Chen et al., 2021a; Zheng et al., 2021a). Different from the methods we introduced above, these works often choose students that are not necessarily simpler than the teacher, which can sometimes outperform the teacher in terms of accuracy, but they could be even slower than the teacher. We only briefly mention them here given they are less related to acceleration. GNN Self-Distillation (GNN-SD) (Chen et al., 2021a) considers teacher-free self-distillation where the knowledge is extracted and transferred between layers of the same GNN. GNN-SD is shown to consistently achieve better accuracy than the standard KD objective $\mathcal{L}_{i}^{KD}$. Combination of Parameterized label propagation and Feature transformation (CPF) (Yang et al., 2021b) combines KD and label propagation (LP) together. The student in CPF includes an MLP module for learning feature information and an LP module for learning structure information. The classification accuracy of a CPF student can generally outperform the accuracy of teacher GNNs after KD. However, since the LP module still has node dependency like the message passing (labels instead of the hidden features are passed as messages), the speed gain is likely limited and not discussed in (Yang et al., 2021b). Cold Brew (Zheng et al., 2021a) utilizes KD to improve GNN performance on nodes with incomplete or missing neighbors, i.e., the cold-start nodes. Cold Brew greatly improves the accuracy on these cold-start nodes, but its inference stage requires a scan through the whole graph to select existing nodes as neighbor candidates for the cold-start nodes. Cold Brew thus has even more node dependency and runs slower than a standard message-passing GNN. More related works include Graph-Free Knowledge Distillation (GFKD) (Deng and Zhang, 2021) transfers knowledge from a teacher GNN to a student GNN via generating fake graphs. GFKD focused on graph-level prediction tasks, for which data instances are independent graphs and the scalability problem caused by node dependence is less urgent.
4.4. Combine Together
Inference acceleration techniques in this section have been combined together to achieve better results. For example, quantization is combined with distillation in (Bahri et al., 2021). In (Bahri et al., 2021), GNNs weights and activations are excessively quantized to binary. This quantization is very ambitious and hard to achieve in one-step. (Bahri et al., 2021) uses a cascaded distillation scheme to gradually quantize more and more parts of the GNN and distill knowledge step-by-step to the next more quantized GNN down the sequence. The final quantized-and-distilled binary GNN can achieve 2X speedup on Raspberry Pi 4B with only moderate accuracy loss.
5. GNN Acceleration: COTS Systems
In addition to the efficient algorithms for GNN training/inference, optimizing the underlying system is essential for improving end-to-end throughputs of GNNs. In particular, existing works accelerate GNN systems from three perspectives: GPU kernel acceleration, user-defined function (UDF) optimization, and scalable training systems. Different from most deep learning (DL) tasks, GNN computation follows a message-passing paradigm where it is highly desirable to have efficient sparse operations; however, this cannot be sufficiently handled by existing DL platforms and infrastructure. Various efficient GNN kernels and UDF optimization techniques have been proposed for bridging the gap between the lack of fundamental support and high demand. Furthermore, due to the irregular dependency among nodes, training a graph with a huge number of nodes and a giant feature matrix is particularly challenging. How to design an efficient scalable GNN training system is still an open research question for the ML system community. Currently, two widely used frameworks for GNNs are PyG (Fey and Lenssen, 2019) and DGL (Wang et al., 2019b).
5.1. GPU Kernel Acceleration
GPUs are powerful hardware accelerators that can greatly speed up computations required in DNN training and inference. Under this context, GPU kernels refer to specialized programs optimized to be executed in parallel on the many processing units of a GPU, enabling DNN computations to be much faster and more efficient. Although GPUs are popular in being used for DL acceleration, accelerating GNNs with GPUs is still challenging due to the unique sparsity and irregularity in graphs. In recent years, many works propose novel GPU kernels for accelerating GNN workloads. Table 3 summarizes recent works targeting different GNN kernels as detailed below.
Table 3. Summary of GPU Kernel Acceleration Methods.
Work Methods Benefits
PCGCN (Tian et al., 2020) Workload Reorder, Dual-Mode Computation Better Data Locality, Coalesced Memory Access
fuseGNN (Chen et al., 2020b) Kernel Fusion, Workload Reorder, Dual-Mode Computation Decreased Memory Traffic, Saved Memory
FusedMM (Rahman et al., 2021) Kernel Fusion Balanced Workload, Saved Memory
TLPGNN (Fu et al., 2022) Workload Reorder, Kernel Fusion, Dynamic Workload Assignment Decreased Memory Traffic, Coalesced Memory Access, Balanced Workload, No Branch Divergence
Zhang et al. (Zhang et al., 2022b) Workload Reorder, Kernel Fusion, Recomputing Embedding Reduced Computation, Reduced I/O, Saved Memory
PCGCN (Tian et al., 2020) improves the data locality in GNN computation by leveraging the unique sparsity pattern in graphs that nodes are usually clustered. As such, processing GNN workloads in terms of subgraphs substantially enhances the data locality. Furthermore, PCGCN employs a dual-mode computation so that SpMM (sparse matrix-matrix multiplication) is used for sparse graphs while GeMM (general matrix-matrix multiplication) is leveraged for dense graphs. For ensuring coalesced memory accesses of features, SpMM processes all edges linking each pair of partitions at each stage. The proposed technique has been implemented on top of TensorFlow. However, PCGCN relies on graph clustering, which may incur non-trivial preprocessing overhead needed for graph partition.
fuseGNN (Chen et al., 2020b) develops two workload abstractions of GNN aggregation, assuming that the input graph adopts the COO format: Gather-ApplyEdge-Scatter (GAS) and Gather-ApplyEdge-Reduce (GAR). Specifically, GAS works with the COO format graph, and processes edge-wise aggregation; GAR handles node-wise aggregation, which is particularly useful for the CSR format as we defined in Section 2.1. Although GAR is usually efficient because it relies on on-chip reduction with a shared output memory, creating CSR from COO is usually time costly, thus degrading the achievable efficiency of GNN computation. As a result, fuseGNN uses GAS for graphs with a low average degree (e.g., Cora(Sen et al., 2008a)) and GAR for graphs with a high average degree (e.g., Reddit (Hamilton et al., 2017a)). All operations in the three kernels, including graph processing for graph format conversion, GAS, and GAR, are fused for alleviating the memory traffic. fuseGNN is developed for PyTorch and can be extended to multi-GPU scenarios.
FusedMM (Rahman et al., 2021) is a CPU kernel but the authors claim that it can be extended to GPU acceleration. The work divides popular graph operators into five types of basic operations that can be launched sequentially (e.g., element-wise additive, reduction operation, element-wise self-operation) and defined by users for designing customized GNN layers. During GNN computations, the operations associated with each target node are fused together for achieving memory reduction by eliminating the necessity of saving intermediate results. Since each CPU thread is assigned multiple target nodes, FusedMM further adopts an edge-balanced partition to ensure a balanced workload among threads. The proposed kernel of FusedMM can be integrated into DGL (Wang et al., 2019b).
TLPGNN (Fu et al., 2022) develops two levels of parallelism: node- and feature-level parallelism. In the first level (i.e., warp-level), TLPGNN assigns each target node to a warp of threads for avoiding branch divergence caused by uneven node degrees. For computing each node’s new embedding, the second level (i.e., thread-level) adopts feature-level parallelism that the threads in the same warp process feature in parallel while processing each edge sequentially, which allows coalesced memory accesses to input features for each source node. To achieve a balanced workload, a dynamic workload assignment is developed, which allocates the next computation task once one hardware resource is released. TLPGNN further performs kernel fusion to avoid unnecessary memory accesses.
Zhang et al. (Zhang et al., 2022b) first decompose a GNN operation into four parts (i.e., scatter across edges, apply edge, gather from edges, and apply node), and then identify three challenges in GNN computations: redundant neural operator computation, inconsistent thread mapping, and excessive intermediate data. Specifically, redundant neural operators can be caused when the same feature transformation is employed to the same node features after they are scattered across edges. To address this issue, the scatter operation can be postponed after the feature transformation is finished. For the second issue, if the thread mappings are inconsistent (e.g., scatter threads are assigned with edge-centric tasks, while gather threads are assigned with node-centric tasks), these threads have to be processed sequentially and cannot be merged. This work leverages unified thread mapping for fusion. The third issue identifies that intermediate results may occupy up to 92% of the total memory requirement, which can be resolved by recomputation during backward propagation.
Overall, workload reordering is necessary for identifying GPU-friendly processing patterns. In particular, most works sequentially compute output nodes and process features in parallel. The former allows kernel fusion and avoids branch divergence while the latter guarantees coalesced memory accesses. Because the graph structure can be highly irregular and can be presented in different formats, some work leverage dual-mode computation and dynamic workload assignment for handling graphs of different properties.
5.2. User-defined Function Optimization
The flexibility in selecting the message, aggregation, and update functions has been a critical component in the success of GNNs under the message-passing paradigm. GNN systems possess the ability for users to natively express these functions using tensor operators in Python and execute them through APIs provided by the system. This programming paradigm is termed user-defined functions (UDFs). Once defined, these functions can be reused throughout the program, increasing code modularity and reducing code duplication. Overall, UDFs provide a powerful tool for creating more efficient and maintainable codes. A few recent works have been developed to ease the deployment of user-defined GNNs.
FeatGraph (Hu et al., 2020c) provides a flexible programming interface on top of the compiler stack for DL systems: TVM (Chen et al., 2018b) for supporting friendly user-defined functions. The users can manually select the tiling factor (the number of features being processed in each thread) for maximizing the cache utilization. However, because TVM lacks the support for sparse operations, FeatGraph only supports SpMM/SDDMM-based operations.
Seastar (Wu et al., 2021) develops a vertex-centric programming interface for user-defined functions, and generates an execution plan that optimizes memory consumption and data locality, resulting in higher performance. Specifically, Seastar first translates the vertex-centric logic into tensor operations via a tracer, and then identifies operator fusion opportunities by the notion of Seastar computation pattern. This approach reduces memory consumption and enhances data locality via fusing operations of each target node by leveraging the merits of vertex-centric programming, similar to FusedMM (Rahman et al., 2021). Additionally, Seastar uses a distributed graph communication library (DGCL) to enable scalable GNN training on distributed GPUs, and its DL backend is MindSpore, an all-scenario AI computing framework (Huawei, 2020).
Graphiler (Xie et al., 2022) develops an automated tool for scatter/gather-related tasks. Specifically, compared with (Zhang et al., 2022b) mentioned in the previous section that manually implements operation reordering for some specific GNN operations, Graphiler enjoys better flexibility by automatically detecting redundant communication and then adjusting their order accordingly, and further fuses all operations after gathering from edges to reduce intermediate memory accesses.
Existing automation tools for GNN computation are not prosperous. This is because concurrent DL backends do not well support sparse operations. We believe this direction is getting mature as several recent backends for sparse DL operations have been proposed. For example, SparseTIR (Ye et al., 2022) is a general compiler infrastructure for sparse DL operations. It develops a set of intermediate representations that can represent operators supported by the aforementioned Seastar, FeatGraph, and Graphiler work.
5.3. Scalable Training Systems
To support the training of GNNs with large-scale graphs, many recent works have developed GNN systems targeting scalable training. From the system perspective, there are two types of scalability:
  • Vertical scalability handles scaling data with restricted computation resources. As the training graph of a GNN can be giant, this direction is highly practical. For the GNN systems pursuing vertical scalability, they usually assume the use of low-cost or unlimited host memory.
  • Horizontal scalability expects better efficiency with scaling resources. However, horizontal scalability is particularly challenging in GNN training because more workers incur more dependency across devices, leading to more significant communication and synchronization overhead and thus degrading the achievable training throughput.
Some concurrent works summarize recent distributed GNN training systems (Shao et al., 2022; Lin et al., 2022). In particular, Shao et al. (2022) classifies recent works in terms of hardware facilities (e.g., multi-GPU, GPU cluster, and CPU cluster). On the other hand, Lin et al. (2022) categorizes all works as full-batch training or mini-batch training. In this work, we focus on scalable GNN training, which is a wider topic than distributed GNN training, as it not only includes efficient distributed training (i.e., horizontal scalability) but also considers large-scale training with limited computing resources (i.e., vertical scalability). Furthermore, we observe that the key contribution of each system overlapped among (between) the categories adopted in existing survey papers, hindering the understanding of their technical differences. To facilitate a clearer comparison of the technical contributions made by different scalable GNN training systems, we categorize them based on the method by which data is loaded onto the computation devices: on-device systems, swap-based systems, and sampling-based systems. An overview of these categories is demonstrated in Figure 4, and we compare their scope in Table 4.
Table 4. Comparison of The Three Types of Scalable GNN Training Systems.
System Type Main Focus Challenges and Contribution
On-device system Full-graph aggregation Communication and memory pattern
Swap-based system Full-graph aggregation Workload scheduler
Sampling-based system Sampling-based aggregation Data movement and caching strategy
Here we briefly summarize these three types of systems:
  • On-device Systems (Figure 4(a)) distributedly store the entire graph and associated features on the computation devices (e.g., GPUs or CPUs). Since each worker only maintains a portion of data, systems based on this storage architecture mostly focus on full-graph aggregation for better leveraging the computation resources. During training, each worker needs to transfer the intermediate embeddings for supporting the full-graph aggregation in the next layer. A naive system that follows this direction suffers from frequent and overwhelming communications as well as exploded memory requirements for caching the embeddings from other workers. Therefore, existing systems of this category mainly focus on maximizing communication and memory savings to enhance horizontal scalability.
  • Swap-based Systems (Figure 4(b)) store all data in (distributed) shared memory and swap it and the intermediate results to workers for computations. This approach enables each node to fully utilize its neighbors’ embeddings. However, developing a workload scheduler that ensures vertical scalability poses two challenges for swap-based systems: 1) balancing workload among different GPUs is non-trivial, and 2) naive full-graph aggregation with limited devices incurs long iteration latency because it only updates the model once per training epoch. Despite these challenges, effective solutions to this problem are crucial for improving the achievable efficiency of these systems and enhancing their overall performance.
  • Sampling-based Systems (Figure 4(c)) also save the whole graph and features in (distributed) shared memory; however, instead of performing full-neighbor aggregation, they adopt neighbor sampling for neighbor aggregation. To implement this approach, each sampler first determines sampled input nodes based on the output nodes for mini-batch training. The features of sampled input nodes are then fetched from the (distributed) shared memory. The primary challenge for sampling-based systems is data movement for mini-batch preparation, which can significantly limit performance. To overcome this bottleneck, these systems focus on developing better data movement and caching strategies to optimize communication. Overall, sampling-based systems enhance both horizontal and vertical scalability, making them a promising solution for GNN training on large-scale graphs.
(a) On-device Systems
(b) Swap-based Systems
(c) Sampling-based Systems
Figure 4. Overview of three types of scalable GNN training systems. The dashed lines in (a) indicate the dependency between nodes from different workers, while the grey nodes in (b) serve only as input nodes for subgraph computations. Model synchronization among workers is not shown for better visual clarity.
Table 5. Summary of Scalable Training Systems.
System Type System Computation Hardware Aggregation Adjust Algorithm Open Source
On-device System CAGNET (Tripathy et al., 2020) GPU Full-graph
Dorylus (Thorpe et al., 2021) Hybrid Full-graph
$P^{3}$ (Gandhi and Iyer, 2021) GPU Sampling
PipeGCN (Wan et al., 2022b) GPU Full-graph
LLCG (Ramezani et al., 2022) GPU Full-graph
BNS-GCN (Wan et al., 2022a) GPU Hybrid
SAR (Mostafa, 2022) CPU Full-graph
Sancus (Peng et al., 2022) GPU Full-graph
Swap-based System NeuGraph (Ma et al., 2019b) GPU Full-graph
Roc (Jia et al., 2020a) GPU Full-graph
GNNAdvisor (Wang et al., 2021a) GPU Full-graph
GNNAutoScale (Fey et al., 2021) GPU Full-graph
GraphFM (Yu et al., 2022) GPU Full-graph
Sampling-based System AliGraph (Zhu et al., 2019) CPU Sampling
PaGraph (Lin et al., 2020) GPU Sampling
DistDGL (Zheng et al., 2020a) CPU Sampling
DistDGLv2 (Zheng et al., 2021b) GPU Sampling
SALIENT (Kaler et al., 2022) GPU Sampling
GNNLab (Yang et al., 2022) GPU Sampling
BGL (Liu et al., 2021a) GPU Sampling
GNS (Dong et al., 2021) GPU Sampling
MariusGNN (Waleffe et al., 2022) GPU Sampling
5.3.1. On-device Systems
One convincing approach for GNN training with large-scale graphs is to partition the data into multiple workers. All workers update the local model using the locally labeled nodes and then synchronize the local models. This design is very similar to distributed data parallelism in DNN training. However, there are two fundamental differences: 1) GNN training requires feature transfer for each layer, which can explode the required communication volume; and 2) the node feature/embedding tensors are substantially larger than the model. These unique communication and memory challenges motivate the development of various GNN training systems. Since systems under this paradigm update models with dependent partitions instead of independent data, this approach is also named as partition parallelism.
CAGNET (Tripathy et al., 2020) develops four communication patterns by partitioning the adjacency matrix of target GNNs into 1D, 1.5D, 2D and 3D formats and then distributing the workload across workers accordingly. This work further analyzes the communication latency for each design. For ensuring a balanced workload, the random partition is adopted. However, because communication does not consider the sparse pattern, broadcasting all features causes redundant data transfer (i.e., workers receive node features that are not required), leading to both significant communication overhead and memory consumption.
Dorylus (Thorpe et al., 2021) is a serverless platform for saving the cost of training GNN models, enabling affordable, scalable, and accurate GNN training. Because conventional GPU servers are expensive and have restricted memory, Dorylus presents a distributed system with much lower-cost CPU clusters and Lambda threads, a representative of serverless threads. The authors notice that the training time of GNNs is dominated by neighbor propagation, rather than neural network operations. As parallel processing does not effectively enhance the speed of neighbor propagation, Dorylus executes 1) the neighbor propagation in CPU clusters for achieving the best performance per dollar, and 2) the neural network operations in Lambda threads for avoiding the necessity of unneeded resources (e.g., storage) that conventional distributed platforms (e.g., CPU clusters) require. To further minimize the network latency between the aforementioned CPU clusters and Lambda threads, Dorylus adopts fine-grained computation and asynchronous execution through pipelined computation and communication including embedding transfer and model synchronization. Following VR-GCN (see Section 3.2), the convergence analysis based on historical embeddings is provided. Overall, Dorylus achieves the best performance per dollar against its baseline methods.
$\mathbf{P^{3}}$ (Gandhi and Iyer, 2021) is a distributed system with sampling-based aggregation focusing on the scenario where a mini-batch can not fit into the memory in a single GPU. Since the main contribution of $P^{3}$ is a novel communication pattern among workers, we place it in the category of on-device systems. Specifically, $P^{3}$ assumes that a GNN model’s hidden embedding dimension is significantly smaller than that of the input features. Based on this assumption, $P^{3}$ leverages inter-layer model parallelism for the first layer (i.e., distributedly computing ${\bm{X}}{\bm{W}}^{(1)}$ by splitting ${\bm{X}}$ along the column dimension) to distribute the burdensome storage of input features and then uses data parallelism (i.e., distributively storing ${\bm{A}}$ for computing node embeddings) for reducing the required communications for transferring intermediate embeddings. Additionally, $P^{3}$ further pipelines communication and computation across mini-batches for hiding the communication overhead. The limitation of this design is that $P^{3}$ degrades the performance when the hidden dimension is increased.
LLCG (Ramezani et al., 2022) is the acronym of “learn locally correct globally” which reduces the communication frequency of model synchronization through periodic data transfer. In particular, each training round of LLCG consists of two phases. In the first phase, each worker constructs local mini-batches and updates local parameters following conventional sampling-based training without data transfer. For the second phase, the model parameters in all workers are averaged in a server that keeps on updating the model through mini-batch training with full-graph aggregation. As a variant of partition parallelism, LLCG does not transfer node embedding but synchronizes model parameters instead. The theoretical analysis shows that the stochastic gradient computed by LLCG is bounded by a constant that tends to zero when increasing the size of the sampling size. Since the achievable accuracy of LLCG mainly benefits from the full-graph aggregation on the server side, we regard it as a type of full-graph aggregation method.
PipeGCN (Wan et al., 2022b) optimizes the straightforward implementation of partition parallelism through pipelining embedding computation and the dependent communication across training iterations. In particular, for distributed full-batch training, each worker needs to wait for dependent neighbors’ embedding (or embedding gradient) to be transferred before performing local forward (or backward) computation. In PipeGCN, each worker no longer needs to wait for the dependent neighbors. Instead, they use the transferred embeddings/embedding gradients in the previous iteration. This design breaks this computation-communication dependency and allows the training system to perform computation and communication in parallel. Although the model update does not follow traditional gradient descent because it introduces stale embeddings and stale embedding gradients, the convergence analysis shows that PipeGCN enjoys the convergence rate of $\mathcal{O}(T^{-\frac{2}{3}})$, with $T$ being the number of training iterations, which is better than popular sampling-based methods ($\mathcal{O}(T^{-\frac{1}{2}})$). PipeGCN further employs embedding/embedding gradient momentum for mitigating the error caused by the stale data transfer.
BNS-GCN (Wan et al., 2022a) accelerates partition parallelism from a different angle. The authors observe three major issues associated with partition parallelism: significant communication overhead, scaled-out memory requirement, and imbalanced memory distribution. They further observe that both communication volume and memory consumption are correlated to the number of boundary nodes (i.e., the dependent neighbors that need to be received for each partition). To mitigate the issues caused by boundary nodes, BNS-GCN simply adopts Boundary Node Sampling, achieving drastically saved communication, reduced memory, and balanced computation, while maintaining full-graph accuracy. The sampling rate can be as low as 0.1 as suggested by the authors.
SAR (Mostafa, 2022) also discovers the exploded memory issue in partition parallelism and proposes Sequential Aggregation and Rematerialization scheme. Specifically, each worker receives dependent neighbors through point-to-point communication sequentially. Once the dependent data from one remote worker is received, the receiver worker consumes it, deletes the data to release the memory, and receives the node embeddings from the next worker. For supporting backward propagation, SAR uses reversible neural networks to efficiently reconstruct (rematerialize) the deleted data.
Sancus (Peng et al., 2022) improves CAGNET by performing expensive broadcast communication selectively. To avoid overwhelming broadcast communication, which may occupy over 80% of the total training time, each worker maintains a copy of other workers’ embeddings and only refreshes them when the embeddings are too stale. Based on this design, Sancus is able to skip considerable rounds of communication, thus achieving better throughput. Three staleness metrics are proposed for the staleness check.
5.3.2. Swap-based Systems.
Because full-graph aggregation with on-device storage has expensive storage requirements for scaling graphs, another direction of full-graph aggregation is to store the entire graph in the external shared memory, which motivates swap-based systems.
NeuGraph (Ma et al., 2019b) and Roc (Jia et al., 2020a) are pioneering works for GNN training with scaling graphs. Given restricted GPU resources, they process output nodes sequentially and store all intermediate results in the host CPU DRAM for the computation of backward propagation. Both works support GNN training with multiple GPUs. NeuGraph only supports distributed training in a single machine, and adopts equal-vertex partition for heuristically balancing workload. Roc strengthens the scalability by supporting multi-machine training. It further leverages a learnable cost model for partitioning a graph toward better workload balance.
GNNAdvisor (Wang et al., 2021a) is another work that targets scalable GNN training with limited devices. Compared with NeuGraph and Roc, GNNAdvisor further explores feature-level partition. It implements a low-overhead cost model to determine the best graph-level and feature-level partition. The current implementation of GNNAdvisor only supports single-device training.
GNNAutoScale (Fey et al., 2021) and GraphFM (Yu et al., 2022) develop scalable algorithms for efficient GNN systems. Specifically, to enable the training within limited GPU memory, GNNAutoScale leverages ClusterGCN (described in Section 3.2). During the training process, all nodes maintain their features and the latest intermediate embedding in the host CPU memory. For better approximating neighbor aggregation during the training, GNNAutoScale accesses all maintained historical embeddings for unsampled neighbors. With this design, all nodes can access all neighbor embeddings although the embeddings of unsampled nodes are stale. GraphFM is a follow-up work of GNNAutoScale that develops GraphFM-IB and improves the approximated neighbor aggregation by integrating feature momentum for alleviating incurred staleness. Their convergence analysis shows that GraphFM-IB does not rely on large sampling size. To further mitigate staleness, a variant named GraphFM-OM is developed by pushing the intermediate embeddings of sampled nodes to update the unsampled nodes, achieving better accuracies.
5.3.3. Sampling-based Systems
The last paradigm uses neighbor sampling for saving both training computation and memory requirement. For reducing data movement, various sampling and caching techniques are proposed.
AliGraph (also known as Graph-Learn) (Zhu et al., 2019) is a CPU-based distributed training framework for a wide range of GNN variants and training algorithms. For supporting the efficient training of a large-scale graph, AliGraph utilizes distributed storage with cached important nodes and an optimized data sampler for fetching dependent node features. Specifically, the system implements four graph partition algorithms for different tasks (e.g., METIS is used for separating sparse graphs, streaming-style partition works on dynamic graphs with frequent edge change). Since distributed storage increases data access overhead, AliGraph caches frequently visited nodes for avoiding repetitive data transfer. The system further implements three sampling strategies for speeding up GNN training and accelerates the samplers with dynamic sampling weights. AliGraph has been deployed at Alibaba for business.
PaGraph (Lin et al., 2020) is a multi-GPU system for GNN training. For reducing the data transfer between CPU and GPU for neighbor sampling, PaGraph heuristically minimizes data movement by caching nodes with large out-degrees in GPUs. To enable multi-GPU training, PaGraph devices a graph partition algorithm to balance training nodes across partitions, achieving approximately balanced workload. Each partition only maintains nodes within the partition and $L$-hop neighbors for an $L$-layer model for avoiding redundant storage.
DistDGL (Zheng et al., 2020a) and DistDGLv2 (Zheng et al., 2021b) are distributed training frameworks based on DGL (Wang et al., 2019b). DistDGL is developed for CPU clusters where each worker maintains a partitioned subgraph in the host memory with its node features and one-hop neighbors. During the training process, each worker first launches processes for mini-batch sampling which create training batches while fetching features of sampled remote neighbors. METIS is used for balancing training nodes across partitions and reducing edge cuts so that the communication across workers is minimized. To further approximately balance workload, multi-constraint partition is adopted for balancing nodes and edges. For the graphs with learnable node embedding, DistDGL adopts asynchronous update of sparse embedding for overlapping communication and computation while enjoying seldom data access conflicts. DistDGLv2 extends DistDGL to heterogeneous graphs and GPU clusters and deploys asynchronous mini-batch generation for fully utilizing all computation resources. Overall, DistDGLv2 is 2$\times$-3$\times$ faster than DistDGL. Both projects have be integrated with DGL and DistDGLv2’s APIs are compatible with DGL.
SALIENT (Kaler et al., 2022) boosts the training of GNNs with an efficient dataloader. This work identifies that for the conventional implementation of GNN training where neighbor sampling is conducted in CPU and the GNN model is computed in GPUs, batch preparation and data transfer takes around 72% total epoch time, which severely degrades the efficiency of GNN training. In particular, batch preparation contains two stages: 1) sampling and creating message-flow graphs; and 2) fetching feature and labels through tensor slicing. The first stage has a large optimization space because neighbor sampling can be accomplished by multiple design choices and implementation (e.g., mapping node indices between the original graph and the sampled message-flow graph, fusing the operations of neighbor sampling and minibatch construction). SALIENT empirically explores a fast sampler that is microbenchmarked in ogbn-products (Hu et al., 2020b) for each individual hop, which yields 2.5x faster speed than PyG. For accelerating the second stage, SALIENT implements C++ threads for parallelly slicing tensors in shared memory. To further mitigate the overhead of data transfer, SALIENT pipelines data transfer and GPU computation. SALIENT also evaluates neighbor sampling for efficient GNN inference.
GNNLab (Yang et al., 2022) improves PaGraph with sampling minibatches in GPUs. This improvement is not trivial because on-device sampling requires extra memory which restricts cache size thus increases missing rates. In addition, simply caching nodes with large out-degree is not optimal because it ignores the pattern of training nodes. Therefore, some highly-connected nodes are not frequently sampled during the training process. To address this issue, GNNLab first performs serveral rounds of pre-sampling. Most frequently visited nodes are selected as cached nodes. In addition to the pre-sampling based caching policy, GNNLab further enhances cache size through space-sharing design. Some GPUs are selected as samplers and the others are trainers. This strategy enhances data locality for all workers. To balance the execution time of samplers and trainers, GNNLab first determines the number of samplers and trainers based on their execution time in one epoch so that samplers can finishes sampling tasks in shorter time, and then, idle samplers are switched to trainers for fully leverage computation resources.
BGL (Liu et al., 2021a) also observes data movement bottlenecks in sampling-based GNN training systems and develops a dynamic GPU cache by leveraging temporal locality among mini-batches, which can be strengthened by selecting training nodes via BFS sequences. However, a straightforward implementation incurs biased label distribution. To mitigate this issue, BGL generates multiple BFS sequences by randomly selecting BFS roots. Furthermore, for each training epoch, the BFS sequences will be shifted to add randomness for batch generation. Under the multi-GPU setting, BGL ensures no duplicated cached nodes because cached nodes can be shared via NVLink and this design saves GPU memory. BGL further develops a scalable graph partition algorithm with BFS while heuristically ensuring a balanced workload. Finally, BGL optimizes resource allocation of data preprocessing by formulating it as an optimization problem.
GNS (Dong et al., 2021) stands for Global Neighbor Sampling, which develops an efficient and scalable sampling algorithm for reducing the data movement between CPU and GPU. To achieve this, GNS periodically select a subset of all nodes and caching their node features in GPUs. During the neighbor sampling process, if the GPU caches contain sufficient neighbors, GNS only samples neighbors from the caches. Otherwise, extra neighbors outside GPUs are sampled. GNS devices two caching strategies: 1) the nodes are cached with the probability proportional to their degrees so that only a small cache needs to be maintained for a power-law graph; 2) the sampling probability for each node is determined by random walks where more frequently visited nodes are more likely to be sampled.
MariusGNN (Waleffe et al., 2022) is a GNN training system targeting deeper vertical scalability, which is developed based on Marius (Mohoney et al., 2021). The above-mentioned works assume that all data can be saved in the host CPU memory. This work considers a more resource-constraint scenario where CPU memory is limited. For reducing expensive I/O operations, MariusGNN develops a sophisticated dataloader, which enjoys reduced I/O operations and continuous external memory access.
DSP (Cai et al., 2023) accelerates GNNs by addressing two main challenges associated with multi-GPU training: low GPU utilization and high communication costs. To address the first challenge, DSP uses a tailored data layout that stores the graph topology and popular node features in GPU memory and leverages NVLink to share GPU caches among all trainers, which allows for efficient graph sampling with multiple GPUs. Additionally, DSP introduces a collective sampling primitive (CSP) that pushes the sampling tasks to data to reduce communication costs. To address the second challenge, DSP employs a producer-consumer-based pipeline that allows tasks from different mini-batches to run congruently, improving GPU utilization.
Overall, the core challenge towards scalable GNN training systems is data transfer. To address this issue, several methods have been widely adopted, such as caching frequently visited nodes to avoid redundant communication, using asynchronous communication to hide communication overhead, and leveraging stale embeddings to reduce communication volume. However, in addition to reducing data transfer, saving memory and balancing workload are also critical factors to consider when designing scalable GNN training systems. Achieving these goals requires a combination of effective algorithms and system optimizations to ensure efficient and effective training on large-scale graphs.
6. GNN Acceleration: Customized Hardware
The increasing interest in GNNs has led to the development of customized accelerators (FPGA or ASIC) for faster processing. While GNNs share similarities with CNNs in network architecture, the differences in computation complexity and communication patterns make numerous CNN accelerators (Sohrabizadeh et al., 2020; Zhang et al., 2018; Chen et al., 2016) unsuitable for GNNs. Specifically, GNNs require Matrix Multiplication (MM) units and have irregular memory access due to the unstructured nature of graphs. On the other hand, although both the Aggregation and Update stages (the main stages of a typical GNN) can be modeled as MMs, their computation and communication pattern differ. The Aggregation stage handles an ultra-sparse adjacency matrix, while the matrices in the Update unit are dense or have a much lower rate of sparsity. As a result, the use of only dense MM units (Wang et al., 2021b) or the sparse MM (SpMM) units (Srivastava et al., 2020; Song et al., 2022) is inefficient for this application. Furthermore, GNNs work with vectors assigned to each node, unlike traditional graph algorithms that use scalar values, resulting in distinct computation and communication requirements. These distinctions have motivated researchers to design specialized hardware modules to effectively process GNNs.
Abadal et al. (Abadal et al., 2021) concisely review the endeavors undertaken in this domain, offering insightful and informative perspectives. We try to accentuate the challenges and the differences in the employed optimization techniques more clearly. Section 6.1 describes the challenges that the designer would face and the decisions one must make in developing the accelerator. Section 6.2 reviews the most prominent accelerators proposed to date and discusses how each work approached the problem.
6.1. Challenges in Customized Hardware Design
Graphs and GNNs possess specific characteristics that necessitate special attention in the design of customized hardware. Specifically, designing customized accelerators for these structures requires consideration of the following unique features:
Wide range of GNN layers. Various GNN models utilize distinct Aggregation and Update methods, in addition to potentially incorporating graph/subgraph pooling layers and non-linear activation functions. These differences may impact the execution bottlenecks and areas requiring acceleration, resulting in a trade-off between flexibility and performance. This is because the capability to support multiple modes of computation necessitates generality, which may prevent us from getting the optimal performance for each specific case.
Computation/communication pattern and sparsity rate disparity. The computation and communication requirements of different steps in a GNN may vary significantly due to the distinct sparsity rates of the matrices involved and their irregular access pattern. For instance, the adjacency matrix is an ultra-sparse matrix, whereas the weight matrix is dense. The node embeddings in general are dense but if they are generated by activation functions such as ReLU, they may create a sparse matrix, but the sparsity rate is much lower than that of the adjacency matrix. Additionally, an end-to-end GNN application may utilize other computation patterns such as a Multi-Layer Perceptron (MLP), which creates a dense multiplication unit to process the node/graph embeddings. These differences in computation and communication patterns pose a significant challenge to GNN hardware designers, as they must address the following sources of disparity and irregular access patterns:
  • Graph (adjacency matrix) sparsity: The graph is a large and ultra-sparse matrix and poses challenges to the memory design. For instance, a graph such as Orkut (Mislove et al., 2007) comprises over 3 million nodes and more than 223 million edges, yet the density is only $2.5\times 10^{-5}$. The large graph size requires an on-chip and off-chip memory hierarchy for storage, while the sparsity translates to irregular accesses of graph elements that span multiple memory hierarchies.
  • Sparsity introduced by activation operators: The introduction of sparsity by activation operators is a relevant aspect to consider for GNN hardware design. Specifically, activation operators like ReLU can produce sparsity in the node embeddings or features. To enable efficient computation on the node embeddings, it is desirable to address both graph and feature sparsity. This can be achieved by implementing zero elimination or bypassing techniques in the computation units. By doing so, we can accelerate the computation and avoid wasting computation cycles on zero-valued operands.
  • Coordination between modules: The differences in data layout caused by sparsity necessitate the coordination of computation modules. Specifically, the Aggregation step in GNNs often requires sparse matrix-vector multiplication (SpMV) or SPMM modules, which typically shuffle the data layout to enable efficient processing. On the other hand, the Update (also called transformation) step usually works on continuous matrix/vector layouts. Consequently, an accelerator must coordinate the data layout between modules. As mentioned before, in some GNNs, the weights or embeddings in the Update stage may also be sparse. Therefore, to further reduce processing latency, the accelerator must further coordinate the computation pattern and sparsity of the Update module with the SpMV and SpMM modules.
Dynamic input and network structure. Given the dynamic nature of graphs, the architectural decisions for GNNs are highly dependent on the specific characteristics of the input graph and the model’s hyperparameters. The memory and computation requirements of a GNN accelerator are significantly impacted by factors such as the size of the graph, its level of sparsity, and the dimensions of the vectors involved. Therefore, the features and properties of each input graph can be highly distinct from one another, including the following:
  • Different scales and structure of graphs: Real-world graphs exhibit a wide range of node counts, ranging from a few nodes (e.g., 10) (Riesen and Bunke, 2008) to millions (Mislove et al., 2007). This variation in scale results in differences in the computation resources needed to perform GNN computations. Moreover, recent studies have demonstrated that graph size can significantly impact the selection of scheduling and optimization techniques (Sohrabizadeh et al., 2022b; Geng et al., 2020). Additionally, the varying number of neighbors of nodes in the graph introduces a workload imbalance in processing them.
  • Dealing with vectors with various sizes per node: Traditional graph algorithms like BFS, SSSP, and PageRank typically assign a single scalar value to each node. In contrast, GNNs utilize long feature vectors for nodes. This difference in representation affects the memory access pattern and parallelism. The use of long vectors allows for intra-node parallelism and data reuse opportunities when applying the same weight matrix to all node embeddings. A GNN is composed of multiple layers that can operate on vectors of different sizes. This leads to changes in the access pattern and the required degree of parallelism across different layers. These differences create additional opportunities for customization in the hardware design of GNN accelerators.
Given the special features of graphs and GNNs, designers must carefully consider how to address the aforementioned critical factors when designing customized hardware accelerators. These include selecting appropriate data structures and hardware architectures that can handle the large and irregular data structures involved in graph processing, optimizing parallelism and exploiting the available sparsity to achieve high-performance computing, balancing the workload across the processing units, minimizing communication overheads between processing units, and adjusting the flow of data to avoid potential bottlenecks. Additionally, designers must also consider how to optimize the accelerator for the specific GNN model and its corresponding computation and communication patterns. Furthermore, given the wide range of GNN models, each with its unique characteristics and performance requirements, the designer must consider the trade-off between generality (flexibility), scalability, and specialization to achieve optimal performance for a given application.
6.2. Summary of The Existing Customized Accelerators
The accelerators presented in this section employ distinct approaches to address the aforementioned challenges, influenced in part by the neural networks and datasets they target. Such a trade-off can affect the attainable peak performance and the generalizability of the proposed accelerators. Naturally, as the application scope of the work narrows, more opportunities arise for customization of the accelerator, thereby improving its performance, albeit at the expense of flexibility to adapt to new experimental settings. These accelerators are primarily intended for deployment, focusing on inference rather than training. The constantly changing network architecture during training presents challenges for adapting the hardware design or requiring regeneration and/or reconfiguration. Consequently, customized accelerators designed for the training stage are rare (Zeng and Prasanna, 2020; Chen et al., 2021c).
The summarized works are presented in Table 6, which compares them based on five main features: graph size, target GNNs, support for layer customization based on the dynamic structure of the input, the approach to handle the inherent sparsity of the target application, and levels of parallelization employed. The prevalent parallelization schemes used in GNN accelerators include parallelizing the feature dimension for all stages, parallelizing nodes in the Update stage, and pipelining the main stages of each layer for intra-layer parallelization. Some approaches also involve exploiting parallel processing of multiple edges in the Aggregation stage, pipelining the execution of multiple layers (inter-layer), or processing multiple graphs in parallel (batch). In the remainder of this section, the related works are categorized based on the flexibility of their workload. Section 6.3 provides a summary of the notable works that propose accelerators to target multiple GNN algorithms. Conversely, another category of works concentrates mainly on the operations of GCNs, which are one of the most widely used GNN algorithms. These works further customize the microarchitecture for GCNs, and their details are presented in Section 6.4.
Table 6. Customized Accelerators Properties
Work Graph Size Target GNNs Layer Customization Sparsity Support Parallelization
EnGN (Liang et al., 2020b) Large, GCN, GRN Reorganize edges Intra-layer, Node-level
Ultra Gated GCN for aggregation Feature-level, Edge-level
Large R-GCN
GraphSage
Rubik  (Chen et al., 2021c) Small, GIN Pre-process to Intra-layer, Node-level
Large GraphSage reorder the graph Feature-level, Edge-level
HyGCN (Yan et al., 2020a) GCN Window sliding and Intra-layer, Node-level
Large GraphSage shrinkage + sampler as Feature-level, Edge-level
GIN a sparsity eliminator (if feature dimension is
DiffPool tool for aggregation small)
FlowGNN (Sarkar et al., 2022) Small GCN, GIN On-the-fly multicasting Intra-layer, Node-level
Large GAT, PNA for data distribution Feature-level, Edge-level
DGN, VN
BlockGNN  (Zhou et al., 2021a) Large GS-Pool Normal SIMD process Intra-layer, Node-level
GCN, GAT Feature-level, Edge-level
GGCN
DeepBurning -GL (Liang et al., 2020a) Large GCN Degree-aware caching Intra-layer, Node-level
GS-Pool Feature-level, Edge-level
EdgeConv
AWB-GCN (Geng et al., 2020) Load-balancing
for both stages: Inter-layer, Intra-layer
Large GCN distribution smoothing, Node-level, Edge-level
remote switching,
evil row remapping
StreamGCN (Sohrabizadeh et al., 2022b) Pre-process for Inter-layer, Intra-layer
Small GCN aggregation - In-situ Feature-level, Node-level
support for update to Edge-level, Batch
prune zeros on-the-fly
GraphACT (Zeng and Prasanna, 2020) Redundancy reduction Intra-layer, Node-level
Small GCN to reduce the number of Feature-level
edges for aggregation
Zhang et al.  (Zhang et al., 2020b) Pre-process to partition Intra-layer, Node-level
Large GCN and reorder the graph, Feature-level, Edge-level
and redundancy reduction
BoostGCN (Zhang et al., 2021a) 3-D tiling of edges, Intra-layer, Node-level
Large GCN nodes, and features; Feature-level, Edge-level
sort-and-combine unit
G-CoS (Zhang et al., 2021b) Small GCN, GraphSage On-the-fly multicasting Intra-layer, Node-level
Large GAT, GIN for data distribution Feature-level
GCoD (You et al., 2022) GCN, GIN Workload balanced denser, Intra-layer, Node-level
Small GAT light workload sparser Feature-level, Edge-level
I-GCN (Geng et al., 2021) Small GCN On the fly reorder Intra-layer, Node-level
Large GraphSage and islandize the graph Feature-level
6.3. Accelerators for General Workloads
As previously stated, a fundamental decision when designing a customized accelerator is to determine the range of applications it aims to support. In this section, we provide an overview of the works that propose a flexible accelerator capable of targeting multiple GNN algorithms. These works either create a unified architecture that handles all computation stages, as discussed in Section 6.3.1, or develop a specialized engine for the primary computation stages due to the distinct computation and communication patterns, as summarized in Section 6.3.2.
6.3.1. All-Stage Unified Architecture.
The works in this category propose a unified architecture to process multiple stages of computation, despite their division into distinct phases, such as the Aggregation and Update steps. The proposed accelerators in this category develop a singular engine capable of handling these stages, even though they have different requirements in terms of computation and communication. We elaborate on their approaches below:
EnGN (Liang et al., 2020b) targets graph convolution network (GCN) (Kipf and Welling, 2016), relational GCN (R-GCN) (Schlichtkrull et al., 2018), gated GCN (Dauphin et al., 2017), GraphSage (Hamilton et al., 2017a), and graph recurrent network (Salehinejad et al., 2017). To process them, the operations are divided into three stages: feature extraction, aggregation, and update. The authors develop a Neural Graph Processing Unit (NGPU) as a unified architecture for executing them. These stages are processed as pipelined matrix multiplications of the feature matrix, neural weights, and adjacency matrix. The NGPU uses a 32-bit fixed-point $128\times 16$ systolic array as the main computation unit, which maps nodes to different rows and their features to different columns. To perform the aggregation, the authors propose a ring-edge-reduce dataflow that connects processing elements (PEs) in the same column with a ring so that the results can be passed through and added together based on the adjacency matrix. As the adjacency matrix is sparse, the authors reorganize the edges to reduce the number of idle PEs during the execution of aggregation. Additionally, high-degree nodes are cached to decrease the number of off-chip transactions since they are reused across multiple operations. EnGN analyzes the effect of different schedules on the number of operations and develops a dimension-aware stage reordering strategy, which changes the order of execution between the aggregation and feature extraction stage based on the input and output dimensions, to reduce the number of operations. Furthermore, it splits the graphs into 2D tiles to fit them into on-chip resources and develops an analytical formula to decide whether for a given layer, a column-order traversal of the tiles results in fewer I/O operations or their row-order traversal.
Rubik (Chen et al., 2021c) develops an accelerator tailored to train GIN and GraphSage. To enhance graph locality and data reuse, the proposed design leverages a pre-processor that reorders the graph to group nodes with common neighbors. The graph subsequently passes through a hierarchical accelerator architecture, consisting of a PE array interconnected via a 2D-mesh network NoC. Each PE embedded in this architecture features a multiply-and-accumulate (MAC) array. To map the graph and features to the accelerator, the authors presented graph-level and node-level mappings. The former entails mapping a window of nodes to one PE, allowing for task-level parallelism, with PEs working on distinct node windows. The latter involves tiling the dense vector-matrix multiplication for the feature update (aggregation and update) using the MAC array. Rubik incorporates a global buffer for the PE array, two private caches (for data and instruction) in each PE, and register files (RFs) in every MAC to facilitate data reuse. G-CoS (Zhang et al., 2021b) is an automated framework that aims to optimize the performance and efficiency of GNNs by co-searching for the best GNN structures and accelerators. This framework consists of two components: (1) a one-shot co-search algorithm for GNN structures and their matched accelerators, and (2) a generic accelerator search space that can be applied to various GNN structures. The accelerator search space in G-CoS is based on a template comprising multiple sub-accelerators that can handle both sparse and dense matrix multiplications. By configuring the settings of each sub-accelerator, such as tiling sizes, tiling order, and interconnection style, they can be specialized to process different clusters of GNN operators with distinct data sizes and sparsity, which ultimately enhances the hardware efficiency. In addition, G-CoS employs local buffers assigned to the intermediate features, index for sparse features (in COO format), and weights and intermediate outputs for each sub-accelerator. These buffers can be interconnected to allow for data sharing and reuse across different sub-accelerators, reducing the need for costly off-chip memory access. G-CoS is a pioneering effort towards automating the search for both GNN model structure and accelerator design knobs.
GCoD (You et al., 2022) is a co-design framework that addresses the challenge of extreme sparsity in GNN inference by optimizing both the algorithm and hardware accelerator. At the algorithm level, GCoD polarizes the graphs into either denser or sparser local neighborhoods without compromising model accuracy. This polarization results in adjacency matrices with two levels of workload, which greatly improves regularity and ease of acceleration. On the hardware level, GCoD integrates a dedicated two-pronged accelerator comprising a dense branch and a sparse branch processor. The dense branch processor employs a chunk-based micro-architecture with several heterogeneous computing modules to balance the workload across different subgraphs from the polarized adjacency matrix. This balance is achieved by allocating computing resources and bandwidth proportionally based on the data volume and operation size involved in each subgraph’s processing. The sparse branch processor accelerates the remaining sparse workloads, which constitute a small portion of nonzero data, mostly on-chip. To store data in a sparse format, the sparse branch processor uses the CSC format, which significantly reduces the storage overhead of adjacency matrices. GCoD’s sparse branch processor also employs a query-based weight forwarding mechanism that flexibly shares on-chip input data from the dense branch processor to enhance data reuse. The weight forwarding is performed on-demand of the sparse branch to achieve more efficient control, and it accesses around 63% of the data, which drastically reduces the off-chip access cost. Overall, GCoD’s algorithm-hardware co-design framework offers better accuracy-efficiency trade-offs than optimizing each aspect separately. By polarizing the graphs into denser and sparser local neighborhoods and integrating a two-pronged accelerator with a dense and sparse branch processor, GCoD provides a comprehensive solution to tackle the challenge of extreme sparsity in GNN inference.
I-GCN (Geng et al., 2021) proposes a new approach called "islandization" to enhance data locality in GNNs. By identifying clusters of nodes with strong internal connections, called "islands," I-GCN aims to minimize off-chip memory accesses and improve on-chip data reuse. To support islandization, I-GCN’s hardware architecture includes an Island Locator and an Island Consumer. The Island Locator searches the graph from multiple nodes in parallel to locate highly-connected nodes termed as hubs. From these hubs, multiple tailored breadth-first-search engines are used to identify islands. The Island Consumer then combines and aggregates the islands and hubs, with a focus on reusing aggregation results among nodes with shared neighbors. To optimize the aggregation process, a ring-based on-chip network is used to distribute partial results and detect reuse opportunities among processing elements (PEs), and an in-network reduction scheme is employed to minimize network communication latency. Overall, I-GCN’s islandization approach and hardware design achieve better data locality and efficiency in GNNs.
6.3.2. Dedicated Module per Stage.
In contrast to the previous category, the works in this category propose a specialized engine for the primary computation stages of the GNN algorithms. By adopting this approach, each engine can be customized according to the requirements of each stage. It can also facilitate the concurrent execution of different stages.
HyGCN (Yan et al., 2020a) targets GCN, GraphSage, GINConv (Xu et al., 2018a), and DiffPool (Ying et al., 2018b) and treats their operations as two main stages, aggregation, and combination (Update in Eq. 1). Due to the distinctive execution patterns of the aggregation and combination stages, the proposed accelerator employs a separate processing engine for each stage and processes them in a dataflow manner. Specifically, the Aggregation engine utilizes parallel single-instruction multiple-data (SIMD) cores to exploit the intra-vertex parallelism resulting from the long feature vectors assigned to each vertex by the GNN algorithms. If the feature vector length is smaller than the total number of cores, the unassigned cores are allocated to other nodes, thereby enabling edge-level parallelism for this stage. As the graph connections are sparsely distributed, a sparsity eliminator is employed to identify a region in the adjacency matrix of the corresponding graph partition that stores the effective edges, using a window sliding and shrinkage technique. To further reduce the computation complexity, the aggregation operation is performed on a sampled set of neighbors. The sparsity eliminator and sampler also help to avoid fetching feature vectors of the nodes that are not connected to any edge in the current graph partition. Subsequently, the Combination engine treats the computation as a dense matrix-vector multiplication (MVM) and is realized as a group of systolic arrays that can work independently or cooperatively depending on the configuration. Each systolic array processes a small group of nodes, with the feature vectors (weights) of the nodes flowing through the row (column) dimension.
FlowGNN (Sarkar et al., 2022) proposes a generic dataflow architecture for accelerating GNNs that can flexibly support the majority of message-passing GNN algorithms, including GCN, GINConv, PNA (Corso et al., 2020), GAT (Veličković et al., 2017), DGN (Beaini et al., 2021), and GNN with virtual node (VN) (Gilmer et al., 2017). The authors argued that common pre-processing techniques used to exploit data locality are not feasible for real-time applications with millions of input graphs with varied structures. Thus, they directly take the graph in COO format without any pre-processing to reorganize the data. FlowGNN’s architecture is based on the idea that any graph-related functions can be expressed with pair-wise message passing. It divides the operations into three steps: 1) the gather phase, which aggregates messages from neighbors; 2) the node transformation (NT) phase, which applies the update function (or feature transformation function) to messages; and 3) the scatter phase, which constructs messages for the next layer by applying the respective message transformation function (e.g., applying the edge weights) on the node embeddings produced in step two. The computations are divided into these steps to integrate edge embeddings and different aggregation functions. The gather and scatter phases handle per-edge computations and are implemented as a single unit (MP unit) that follows the NT unit in a dataflow fashion. The NT unit handles per-node computations. As such, the MP (NT) unit consists of parallel PEs that parallelize the edges (nodes). Each of the PEs further parallelizes the features. To avoid bank conflicts, each of the PEs in the MP unit processes one particular bank of edges. FlowGNN uses an NT-to-MP adapter to perform on-the-fly multicasting of the node embeddings generated by the NT unit to their right PE in the MP unit.
BlockGNN (Zhou et al., 2021a) aims to compress the update stage (feature transformation) of GNN models for GraphSage-Pool (GS-Pool), GCN, GGCN (Marcheggiani and Titov, 2017), and GAT that require a large weight memory, high computing resources, and high computing latency when the feature size is large. To achieve this, they leverage block-circulant weight matrices, which are configured in the form of block-circulant matrices that can be accelerated by Fast Fourier Transform (FFT) and Inverse Fast Fourier Transform (IFFT). This transformation reduces the computation intensity and required storage for the update stage. The computation engine of BlockGNN is composed of two main units, the CirCore unit for the update stage and the VPU (vector processing unit) for the aggregation stage and non-linear operations. The CirCore unit is a three-stage pipeline, consisting of an FFT unit that converts data from the spatial domain to the spectral domain, a systolic array that performs element-wise product and accumulation, and an IFFT unit that transforms the results back to the spatial domain. The VPU is organized as a SIMD unit with m lanes, each with a parallel factor of 16.
DeepBurning-GL (Liang et al., 2020a) is an automated framework designed to facilitate the development of hardware accelerators for GNNs. The framework takes a GNN model as input, which is designed using software frameworks such as DGL and PyG, and generates a hardware accelerator for the target FPGA platform. For the hardware design, the authors provide three templates. The first template is the GNN computation template that handles dense and sparse multiplications. Dense multiplications are implemented via a systolic array or a dot production array, while sparse multiplications are handled by an array of SIMD units or a ring-reduce topology (similar to EnGN). The second template is the memory template, which consists of a distributed on-chip buffer, cache, or degree-aware hierarchy on-chip buffer to support both regular and irregular memory accesses. The third template is the graph manipulation template, which supports graph sampling and reconstruction in various GNN models. To adapt to the specific characteristics of the target model and FPGA platform, the authors develop analytical models that analyze the processing requirements and adopt a design space exploration to tune hardware parameters for the required templates based on the graph properties, network architecture, and resource constraints. The authors validate their framework on various GNN algorithms, including GCN, GS-Pool, R-GCN, and EdgeConv (Wang et al., 2019a).
6.4. Accelerators for Special Workload
The literature in this field endeavors to direct their efforts toward a more specialized algorithm. More specifically, it concentrates on a singular GNN algorithm (here, mostly GCN), and subsequently tailors the proposed solutions accordingly. Given that the target model may contain multiple layers, the first subgroup develops a deep pipeline with custom layer modifications, which are elaborated upon in Section 6.4.1. Conversely, the second subgroup adopts a fixed hardware approach for all layers, which is described in detail in Section 6.4.2.
6.4.1. Layer-Customizable Deep Pipeline.
The studies presented in this section argue that a fundamental optimization in a GCN accelerator should involve reducing the expenses associated with memory transactions. To this end, they construct a deep pipeline that circumvents the need for off-chip (global) memory transactions concerning intermediate results among distinct GCN layers. Through the allocation of specific engines for each layer, they can additionally tailor hardware parameters based on the individual workload of each layer.
AWB-GCN (Geng et al., 2020) argues that many real-world graphs follow the power-law distribution, which implies that the number of nodes with a given degree x follows a proportional relationship with $x^{-\beta}$ for a constant $\beta$. This leads to a situation where a small number of rows/columns in the adjacency matrix contain the majority of non-zero elements, while the rest contain only a few non-zero elements. The authors also note that ReLU activation leads to an abundant number of zero elements, making it more efficient to treat the feature transformation step (Update step) as a sparse computation. In response to these observations, the authors propose an accelerator architecture called Autotuning-Workload-Balancing (AWB) GCN, which includes distribution smoothing, evil row remapping, and remote switching. Since GCNs have linear aggregation and update functions, the authors propose executing the update stage before the aggregation stage since this approach helps reduce the number of non-zero operations. This is because by this change, in both steps, we are dealing with two SpMM as opposed to an SpMM and a dense matrix multiplication. AWB-GCN’s architecture supports inter-layer parallelism in addition to the commonly used intra-layer parallelism and is based on the inner-product matrix multiplication. The authors partition the graph and implement a task distributor to navigate the data to idle PEs to balance the workload. The architectural optimizations are designed to overcome the load imbalance problem. Distribution smoothing is local and tries to balance the load among neighboring PEs within up to three hops. The remote switching addresses regional clustering by exchanging the workloads of under- and overloaded PEs. Evil row remapping distributes rows with the most non-zeros to the most under-loaded PEs.
StreamGCN (Sohrabizadeh et al., 2022b) presents an efficient and flexible GCN accelerator for streaming small graphs - from the DRAM, the host CPU, and through the network - and exploiting all the available sparsity. The authors argue that in memory-bound applications like this, it is essential to minimize memory transactions, and they propose a scheduling mechanism to achieve this. In this scheduling, input data is fetched from global memory (DRAM) only once and is reused for all corresponding computations before being evicted. The proposed GCN accelerator includes dedicated modules for each layer, which not only allows customization of parallelization based on the workload of each layer but also enables inter-layer pipelining. As a result, the intermediate results are directly passed to the modules for the next layer through on-chip FIFOs which further helps with decreasing the number of DRAM transactions. To avoid read-after-write (RAW) dependencies in the aggregation unit, the edges are pre-processed and reordered. Both the aggregation unit and the feature transformation unit contain multiple SIMD PE with the SIMD dimension parallelizing the node features. The PEs in the aggregation unit process different pairs of neighbors in parallel with their respective edge weight (edge parallelism), while in the feature transformation unit, they parallelize the nodes. These units are connected in a dataflow fashion. Similar to AWB-GCN, StreamGCN treats the feature transformation step as a sparse computation. However, due to the smaller size of the target graphs, the method is based on outer-product-based multiplication. This scheduling is shown to reduce the number of pipeline stalls due to RAW dependencies. Since the intermediate results are to directly be processed as they are being produced, StreamGCN has in-situ sparsity support by developing a mechanism to prune the zero elements of the node embeddings on-the-fly, while they are being generated. This is realized by a pruner and an arbiter unit. The pruner pre-fetches the results at a higher rate than the next layer’s processing rate to pass non-zero elements through FIFOs. The arbiter then checks them for RAW dependencies and dispatches enough elements to fill the PEs. Overall, the proposed StreamGCN architecture is ideal for real-time or near-real-time graph search and similarity computation for many biological, chemical, or pharmaceutical applications.
6.4.2. All-Layer Unified Architecture.
As seen in the previous section, a GCN algorithm consists of multiple layers that have distinct features (e.g., node embedding dimension). In contrast to the works detailed in Section 6.4.1, the studies presented in this category put forth the idea of constructing a more adaptable architecture across distinct layers and utilizing the same engines for all layers. Since the works in this category focus on a specific workload, namely GCN, they offer greater possibilities for customization than those presented in Section 6.3, as indicated below:
GraphACT (Zeng and Prasanna, 2020) develops an accelerator for training GCNs targeting small graphs on CPU-FPGA heterogeneous systems. The small graphs are created by sampling subgraphs of the input graph to be able to fit them into the on-chip resources. GraphACT does not work directly with the normalized adjacency matrix and rather defines three types of multiplications for GCNs: 1) dense (weights) - dense (embeddings), 2) binary sparse (adjacency matrix) - dense, 3) diagonal (degree matrix) - dense. The first multiplication is performed using a 2D systolic array called the Weight Transformation module, where nodes are mapped to the row dimension and output features to the column dimension. The Feature Aggregation module implements steps two and three. The third multiplication (Step three) is treated as scaling the rows of the second matrix in the multiplication with minimal hardware overhead. GraphACT implements a redundancy reduction technique to reduce the computation cost of the second step, which seeks to pre-compute the repeated aggregations on CPU. This is possible since GraphACT directly works with the adjacency matrix in its binary (not normalized) format. The redundancy reduction algorithm identifies the common pairs of neighbors and merges each of the most repeated ones in a node. This is done as a pre-processing step on CPU since it is communication-intensive. The Feature Aggregation module has a 1D accumulator array that parallelizes the feature dimension. It adds the features of the identified node pairs and performs the aggregation based on the adjacency matrix of the merged graph (generated after merging the node pairs). Finally, it scales the node features based on the node degrees (Step 3). Both the Feature Aggregation and Weight Transformation modules are reused for all the $L$ layers of the GCN.
Zhang et al. (Zhang et al., 2020b) focus on accelerating GCN inference for large graphs. The approach begins by partitioning the input data into smaller sizes that can fit into the on-chip resources for processing. Following this, a redundancy reduction scheme, such as GraphACT, and a node-reordering phase are applied to decrease global memory access. Upon completion of the pre-processing stage, the graph is passed to a hardware architecture that incorporates two primary modules: Aggregation and Weight Transformation. The Aggregation module employs parallel vector accumulators (VAs) for sparse matrix multiplication in the GCN aggregation process. Here, each VA parallelizes the feature dimension and different VAs process distinct edges. The Weight Transformation module leverages a systolic array to implement the update stage. These modules are bi-directionally connected to enable different modes of scheduling, where the order of executing the aggregation and update stage can vary based on the GCN and dataset properties.
BoostGCN (Zhang et al., 2021a) presents a framework aimed at optimizing GCN inference on FPGAs. The hardware architecture of BoostGCN includes two primary modules, namely the feature aggregation module (FAM) and the feature update module (FUM). These modules are interconnected through internal buffers that cache intermediate results, and the connection between them is bi-directional, as in Zhang et al.’s work. To reduce memory traffic and address workload imbalance, the authors propose Partition-Centric Feature Aggregation (PCFA), which leverages tiling in three dimensions: edges, nodes, and features. FAM operates on multiple edges in parallel and further parallelizes the feature dimension for them. Additionally, FAM uses a sort-and-combine (SaC) unit to sort the inputs and combine them if two of them have the same destination, which reduces congestion. BoostGCN employs two types of FUM based on the sparsity of the feature matrix (node embeddings) – Sparse-FUM and Dense-FUM. Sparse-FUM uses the same structure as FAM, with a format transformation module in the beginning that transforms the input to COO format. Dense-FUM, on the other hand, implements a 2D systolic array.
Table 7. Performance and Validation Methodology of the Customized Accelerators
Work Validation Methodology End-to-end Measurement Reported Speedup
EnGN (Liang et al., 2020b) ASIC (TSMC 14nm with HBM 2.0 at 1GHz) 1802.9×1802.9$\times$1802.9 ×, 19.75×19.75$\times$19.75 ×, and 2.97×2.97$\times$2.97 × over CPU
cycle-accurate simulator (PyG &\&& DGL), GPU (PyG &\&& DGL), and HyGCN
Rubik  (Chen et al., 2021c) ASIC (45nm at 500MHz) 3.4246.7×3.42-46.7$\times$3.42 - 46.7 ×, and 1.314.16×1.3-14.16$\times$1.3 - 14.16 × over
cycle-accurate simulator PyG-GPU, and Eyeriss (Chen et al., 2016)-like accelerator
HyGCN (Yan et al., 2020a) ASIC (TSMC 12nm with HBM at 1GHz) 1509×1509$\times$1509 ×, and 6.5×6.5$\times$6.5 × over
cycle-accurate simulator PyG-CPU, and PyG-GPU
FlowGNN (Sarkar et al., 2022) FPGA (Xilinx Alveo U50 at 300MHz) 24254×24-254$\times$24 - 254 ×, 1.3477×1.3-477$\times$1.3 - 477 ×, and 1.26×1.26$\times$1.26 × over
on-board evaluation PyG-CPU, PyG-GPU, and I-GCN
BlockGNN  (Zhou et al., 2021a) FPGA (Xilinx ZC706 at 100MHz) 2.3×2.3$\times$2.3 ×, and 4.28.3×4.2-8.3$\times$4.2 - 8.3 × over
on-board evaluation and analytical modeling CPU (TensorFlow-based), and HyGCN
DeepBurning -GL (Liang et al., 2020a) FPGA (Xilinx ZC706, KCU1500, and Alveo U50) 7.43346.98×7.43-346.98$\times$7.43 - 346.98 ×, 3.716.5×3.7-16.5$\times$3.7 - 16.5 ×, and 6.28×6.28$\times$6.28 × over
on-board evaluation at 100, 200, and 200MHz DGL-CPU, DGL-GPU, and HyGCN on FPGA
AWB-GCN (Geng et al., 2020) FPGA (Intel Stratix 10 SX) 3255×3255$\times$3255 ×, 80.3×80.3$\times$80.3 ×, 5.1×5.1$\times$5.1 × over
on-board evaluation PyG-CPU, PyG-GPU, and HyGCN
StreamGCN (Sohrabizadeh et al., 2022b) FPGA (Xilinx KU15P, Alveo U50 and U280) $8.2\times$ 18.2 ×, and 12.126.9×12.1-26.9$\times$12.1 - 26.9 × over
on-board evaluation at 201, 279, 290MHz PyG-CPU, and PyG-GPU
GraphACT (Zeng and Prasanna, 2020) FPGA (Xilinx Alveo U200 at 200MHz) (training speed) 1215×12-15$\times$12 - 15 ×, and 1.11.5×1.1-1.5$\times$1.1 - 1.5 × over
on-board evaluation CPU, and GPU
Zhang et al.  (Zhang et al., 2020b) FPGA (Xilinx Alveo U200 at 250MHz) 30×30$\times$30 ×, and 2×2$\times$2 × over
on-board evaluation CPU, and GPU - TensorFlow and C++ (Cuda)
BoostGCN (Zhang et al., 2021a) FPGA (Intel Stratix 10 GX 10M at 250MHz) 100×100$\times$100 ×, 30×30$\times$30 ×, $3$-45×45$\times$45 × over CPU (PyG &\&& DGL),
on-board evaluation GPU (PyG &\&& DGL), and FPGA ((Zhang et al., 2020b) &\&& HyGCN)
G-CoS  (Zhang et al., 2021b) FPGA (Xilinx VCU128 at 330 MHz) on-board evaluation 5.52×5.52$\times$5.52 ×, and 1.92×1.92$\times$1.92 × speedups over
HyGCN, AWB-GCN
GCoD  (You et al., 2022) FPGA (Xilinx VCU128 at 330 MHz) on-board evaluation 7.8×7.8$\times$7.8 ×, and 2.5×2.5$\times$2.5 × over HyGCN, AWB-GCN
I-GCN  (Geng et al., 2021) FPGA (Intel Stratix 10 SX at 330 MHz) on-board evaluation 368×368$\times$368 ×, 453×453$\times$453 ×, 16×16$\times$16 ×, and 5.7×5.7$\times$5.7 × over PyG-GPU, DGL-GPU, SIGMA, and AWB-GCN
Table 7 shows a review of accelerators, focusing on their validation methodology and reported speedup. These accelerators have been implemented on FPGAs or simulated as an ASIC design. A prevalent baseline for reporting the accelerator’s performance is CPU and GPU, with most works utilizing the state-of-the-art implementation of PyG for this purpose. The reported results may vary regarding whether they exclusively measure the runtime of the execution of the main kernel for their accelerator or include other associated overheads and present an end-to-end runtime.
7. GNN Acceleration: Special Graphs and Special GNNs
In the sections above we talked about acceleration techniques for the regular message-passing GNNs on regular graphs. However, many practical problems are better modeled with special graph types, e.g., heterogeneous graphs and dynamic graphs (both defined below), and solved with more sophisticated GNN architectures, e.g., heterogeneous GNNs and GNNs with attention. Many methods we discussed above may be applied to these special cases if their specialties are ignored, e.g, treating a heterogeneous graph as if it is homogeneous. However, the result would be suboptimal such as having high accuracy loss. In this section, we discuss techniques for handling these special cases especially.
7.1. GNN Acceleration on Heterogeneous Graphs
The graphs we talked about so far have nodes/edges of the same type, and they are called homogeneous graphs. The complex graphs in the real world are often associated with various types of nodes and edges (Dong et al., 2020; Javari et al., 2020; Huang et al., 2022), establishing a heterogeneous structure. For example in a citation graph, nodes can be authors, venues, and papers, and edges can denote an author published a paper, a paper published in a venue, etc. Different types of nodes and edges have substantially different meanings from each other and it is better to distinguish them explicitly. Heterogeneous graphs as a new data structure are defined below.
Definition 7.1 ().
A heterogeneous graph is defined as a directed graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ associated with a node type mapping function $\phi:{\mathcal{V}}\rightarrow{\mathcal{A}}$ and an edge type mapping function $\tau:{\mathcal{E}}\rightarrow{\mathcal{R}}$. Each node $v\in{\mathcal{V}}$ belongs to one node type $\phi(v)\in{\mathcal{A}}$ and each edge $e\in{\mathcal{E}}$ belongs to one edge type $\tau(e)\in{\mathcal{R}}$.
Many GNNs are specially designed for heterogeneous graphs (Hu et al., 2020a; Zhang et al., 2019b; Long et al., 2021). They face similar scalability challenges as regular GNNs on homogeneous graphs and require acceleration. In the previous sections, we discussed acceleration methods that could be used for GNNs on heterogeneous graphs by disregarding their heterogeneity (i.e., different node and edge types). However, applying acceleration methods in a homogeneous manner by ignoring heterogeneity often leads to a significant decrease in performance. For example, for sampling algorithms, if a message passing step sample node neighbors uniformly without considering the node types, some infrequent types are likely missed out. Completely missing one type of node can downgrade the GNN performance a lot (Zhang et al., 2019b). A better sampling algorithm on heterogeneous graphs should perform a weighted sampling to include neighbor nodes of all types. Similarly for model pruning, heterogeneous GNNs often have separate model weights in the update function for each edge types (Schlichtkrull et al., 2018), and ignoring heterogeneity may prune the weights completely for some edge types, which makes the GNN unable to pass messages that type of edges and downgrade the performance. Therefore, accelerating GNNs while taking into account heterogeneity presents additional challenges and requires special treatment to capture heterogeneity. In the following, we will provide a summary of recent research on modeling and accelerating GNNs on heterogeneous graphs.
HetGNN (Zhang et al., 2019b) is a GNN on heterogeneous graphs that introduces a novel heterogeneous neighbor sampler. The sampler belongs to the node-wise sampling category discussed in Section 3.2.1 but overcomes a special challenge on heterogeneous graphs. On heterogeneous graphs, node types in a target node’s receptive field can be very imbalanced, e.g., some types are very infrequent compared to other types. HetGNN argues that a regular sampler may miss some less common node types completely and result in insufficient node representations. Therefore, instead of randomly sampling, the HetGNN sampler first runs random walks with restart (RWR) in the neighborhood and then chooses nodes to form the computation graph according to their hit frequency by RWR. The sampler will also balance node types, such that the top frequent nodes of all types are chosen. With this RWR-based and type-balanced sampler, HetGNN can achieve fast training with good accuracy performance.
HGT (Hu et al., 2020a) is a transformer-based GNN on heterogeneous graphs, which also introduces an efficient sampler to make the model handle Web-scale graph data with billions of edges. The HGT samples subgraphs in mini-batches. The sampler is able to 1) maintain a similar number of nodes and edges for each type in each subgraph and 2) keep the sampled subgraph dense. Specifically, for each node type it maintains a node budget and iteratively samples the same number of nodes with an importance sampling strategy to reduce variance. The sampling probability for each node type is calculated based on the square of the cumulative normalized degree of other nodes that have already been sampled in its budget. The model is shown to achieve state-of-the-art results on the Open Academic Graph (OAG) which contains millions of nodes and billions of edges.
Pigeon (Wu et al., 2023) is an intermediate representation (IR) and code generator for end-to-end training and inference of relational graph neural networks (RGNNs) (Schlichtkrull et al., 2018) on heterogeneous graphs. Pigeon addresses the performance challenges posed by RGNNs by decoupling the model semantics, data layout, and operator-specific schedule, and expressing these opportunities to allow them to be integrated into the design space as integral elements. The proposed solution includes a compact tensor materialization scheme and a linear operator fusion pass to achieve both training and inference. The compact tensor materialization scheme reduces the resources spent on computing and storing common subexpressions in RGNNs. Specifically, certain edge data are determined by source node features and edge types, and instead of computing and storing such data for each edge, the system computes and stores the data once for each (edge type, unique node index) pair. This reduces the amount of computation and storage required for common subexpressions. The linear operator fusion pass switches the order of linear operators, such as linear layers and inner products. When a linear operator is followed by another linear operator, their order may be switched to reduce the strength.
7.2. GNN Acceleration on Dynamic Graphs
Spatio-temporal data appears frequently in scientific research and real-world applications (Huang et al., 2020; Sanchez-Gonzalez et al., 2020). Dynamic graphs have shown great performance for modeling such spatio-temporal data to solve problems like molecular dynamics in biology (Yang et al., 2021a), information diffusion over social networks (Wang et al., 2021c), and epidemic spread in public health (Huang et al., 2021a). Dynamic graphs are graphs that can change over time. Changes include the addition and removal of nodes and edges to capture the evolving nature of the data. The evolving nature has imposed additional challenges for corresponding learning and reasoning on dynamic graphs as opposed to static graphs. Therefore, researchers have developed special spatio-temporal GNNs to jointly model the graph structural information on the spatial aspect and the evolving dynamics on the temporal aspect. The regular message passing in Equation (1) has been modified as is shown in Equation (18) for spatio-temporal GNNs, where the node representations are updated by both their spatial neighbors and their historical states.
(18) ${\bm{h}}_{v}^{t}=\textsc{UPDATE}({\bm{h}}_{v}^{t-k:t-1},\textsc{AGGR}(\{\phi({\bm{h}}_{v}^{t-1},{\bm{h}}_{u}^{t-1})|(u^{t-1},v^{t-1}\in{\mathcal{E}}^{t-1}\}))$
Here ${\mathcal{E}}^{t-1}$ denotes the edges for graph at timestamp $t-1$, and sometimes the edges can be latent and need to be inferred first before the message passing.
Spatio-temporal GNNs are often applied to computation-heavy problems like the simulation of large-scale dynamical systems (Sanchez-Gonzalez et al., 2020), which have similar scalability concerns as regular GNNs on static graphs and require acceleration. However, spatio-temporal GNN acceleration incurs special challenges because model training needs to be done chronologically and graph operations (modification and sampling) need to consider both spatial and temporal neighbors. For example, the sampling methods introduced in Section 3.2 may be individually applied to dynamic graphs at each timestamp $t$, but they can fall short in performance as they only consider spatial properties of nodes but omit the temporal influence. In this section, we summarize recent works on spatio-temporal GNN acceleration, which are often applied for learning and reasoning over large-scale dynamic graphs.
TGL  (Zhou et al., 2022) proposes a unified framework for accelerating continuous-time temporal GNN training. It designs a temporal-CSR data structure for rapid access of temporal neighborhood candidates, and a parallel sampler is further equipped to accelerate temporal neighborhood sampling. Also, it proposes a novel random chunk scheduling technique for training acceleration where simply increasing the batch size would result in the obsolete node memory issue. TGL  (Zhou et al., 2022) shows its great capability in handling large-scale dynamic graphs where it achieves an average of 173$\times$ speedup on a multi-core CPU compared with the baselines.
MTGNN (Wu et al., 2020a) applies GNNs to model the spatio-temporal graph for solving the multivariate time series prediction problem, where variables are dependent on each other but their dependence (graph structure) is latent. Naively assuming a fully-connected graph structure will result in $N^{2}$ edges and impose huge GNN computation time. MTGNN proposes a graph learning module to automatically extract the uni-directed relations among nodes, which greatly simplifies the structure and accelerates GNN computation. MTGNN also applies clustering to reduce memory usage by the intermediate states of nodes and save memory access time. Specifically, the edge search space is reduced by first randomly separating nodes into clusters and then only forming edges within each cluster. The clustering process for nodes is repeated in each iteration, which allows every pair of nodes in the same cluster to be possibly connected.
GraphODEs (Huang et al., 2020, 2021a) combine the expressive power of GNNs with the principled ordinary differential equation (ODE) modeling for learning dynamical systems. LG-ODE (Huang et al., 2020) is a latent ODE generative model, which models multi-agent system dynamics with fixed graph structure in a continuous manner. The ODE function in LG-ODE is a GNN to capture the continuous interaction among agents, and the latent initial states for ODEs are inferred through another GNN-based encoder. Towards acceleration, existing discrete neural models learn a fixed-step transition function that takes the system state at time $t$ as input to predict the state at time $t+1$. An adaptive ODE solver will automatically adjust the times, where the graph interaction module is called to provide the balance between model performance and time consumption. Additionally, LG-ODE is especially superior at making long-range predictions. It is able to adjust its prediction length without re-feeding its previous predictions at the same fixed step during training, which greatly accelerates model inference. Moreover, since LG-ODE is only capable of handling dynamic graphs with evolving nodes but fixed edges, the follow-up work CG-ODE (Huang et al., 2021a) is proposed to jointly model the evolution of nodes and edges through two coupled ODE functions.
GN-ETA (Derrow-Pinion et al., 2021) proposes a large-scale GNN model for estimating the time of arrival (ETA) in real time. Supersegment-based acceleration is applied so that the complex interaction of traffic roads and the traffic conditions evolution can be joint modeled over time. In particular, GN-ETA segments the traffic network into several supersegments and is trained to predict both the across-supersegments time and within-supersegments time. For model inference, the sequence of supersegments will be put in a route sequentially where the later supersegements would utilize the prediction time of an earlier one to determine its relevant prediction horizons for scaling up. Additionally, GN-ETA adopts the MetaGradients methodology (Xu et al., 2018b) to stabilize the training process raised by uneven query batches that are commonly seen in real-life applications. GN-ETA is able to handle web-scale applications like Google Maps. Recently, the follow-up work CompactETA (Fu et al., 2020) proposes to encode temporal information using positional encoding instead of recurrent networks to further accelerate the model.
Overall, spatio-temporal GNNs for dynamic graphs are extensions of static GNNs with special treatment for the temporal aspect. Traditional methods usually employ recurrent neural networks to capture the evolving nature of networks over time, but are usually time-consuming and fail to capture the long-term dependency. Recent advances to use positional encoding and neural ODEs, etc have shown promising results in speeding up dynamic GNNs. The works discussed above accelerate spatio-temporal GNNs in various ways to scale up them to computation-heavy dynamic system modeling and web-scale applications.
The proposed hardware accelerators in Section 6.2 have demonstrated the capability to support various graph topologies, provided that the graph is within the scale specified by the accelerator. This is attributed to the implementation of preprocessing techniques, such as graph re-ordering and tiling, within these works. However, challenges arise in supporting dynamic and heterogeneous graphs when the algorithm utilized to process them is altered. To address this issue without necessitating a complete redesign of the accelerator for each algorithmic change, software-hardware co-design can be extended. On the software front, an analyzer can be developed to evaluate the requirements of the graph and target algorithm and determine the optimal partitioning, mapping, and scheduling of operations to enable efficient execution on the hardware. On the hardware side, options include moving towards more general design architectures or supporting multiple specialized architectures that the software component can select based on the application requirements.
8. Future Directions
While significant progress has been achieved, the investigation into GNN acceleration remains in its early stages and has not kept pace with the rapid expansion of graph data. Numerous promising research avenues exist within this field, particularly in terms of acceleration from a COTS system and customized hardware perspective, which have only recently begun to garner attention. Additionally, there is a need to explore acceleration techniques tailored for heterogeneous and dynamic graphs. This section highlights several captivating research directions that warrant further exploration.
Better Graph Modification Algorithms Improving the efficiency of graph modification methods, either coarsening, sparsification, or condensation, represents a promising avenue for future research. In many graph datasets, nodes often have an excessive number of neighbors, which may provide redundant information or even noise. To accelerate GNN training, it may be possible to remove these neighbors more aggressively at the pre-processing stage. For example, through sparsification without preserving the global graph properties as practical GNNs are trained for preserving locality. Graph condensation is another promising technique that can enhance training efficiency, but it currently requires running the training process at least once to identify the best condensed graph. If the optimal graph can be determined earlier, e.g., identify a good condensed graph by training only a few epochs, this technique would become much more useful. Good graph modification methods can also benefit the automating designing and tuning GNN architectures. In practice, preprocessed simpler graphs will significantly accelerate the whole pipeline of GNN development.
Better Sampling Algorithms There are still several potential improvements for future research in sampling methods for GNNs. One potential direction is to investigate hierarchical sampling methods that can exploit the multi-scale structure of graphs to improve sampling efficiency and capture multi-scale features of graphs. Hierarchical sampling can be particularly useful for large-scale graphs with complex structures, such as social networks, where nodes can be organized into clusters or communities at different levels. Another possibility is to further investigate the sampling method for dynamic graphs, where nodes and edges change over time. Individually applying sampling to dynamic graphs at each timestamp is still not very efficient and fails to capture the temporal dynamics of nodes and edges on dynamic graphs.
Combined Inference Acceleration Algorithms Inference acceleration methods, such as pruning, quantization, and distillation, are often orthogonal to each other and can be combined to further improve the speed. Some initial steps have already been taken, such as leveraging distillation when ambitiously quantizing weights to binary. As the inference stage is more flexible training, e.g., no concerns like training stability and model convergence, these methods can be more easily combined in various ways. For instance, pruning and quantization can be combined to achieve a sparse model with low precision. Additionally, pruning can be used in conjunction with distillation to further refine the distilled student model.
COTS Systems Existing work mainly focuses on improving the run-time efficiency of GNNs. In contrast, the memory requirement can often be the bottleneck of accelerating GNNs on large-scale data, which has not been well studied yet. Various techniques including rematerialization, quantization, and data pruning can be considered for reducing the memory overhead. In addition, because most GNN models have irregular computation and data access patterns, the advanced training/inference algorithms can be system unfriendly. This makes system-algorithm co-design for system-friendly GNN algorithms a promising research direction. Finally, full-stack development for GNN systems is still missing in the existing literature. Combining existing techniques at different levels (e.g., algorithm, system, hardware) towards an ultra-efficient GNN system is another interesting research topic.
Customized Hardware As explained in Section 6, memory access presents a significant bottleneck in GNN acceleration. Thus, a key aspect of accelerating the processing of GNNs is to minimize the amount of data that must be fetched from off-chip memory. In addition to the techniques employed by the works reviewed, such as data reuse, on-chip memory caching, and dataflow optimization, exploring approaches such as pruning, quantization, and compression may reduce the volume of data fetched from off-chip memory. Furthermore, a critical requirement is to develop a more advanced automated design generator that facilitates GNN accelerator design. Techniques such as Neural Architecture Search (NAS)(Lin et al., 2021), learning-based models for accelerated hardware design evaluation(Sohrabizadeh et al., 2022a), and intelligent design space exploration to navigate through design candidates(Sohrabizadeh et al., 2022c; Mirhoseini et al., 2021) can be suitably adapted for GNN accelerator design.
9. Conclusion
In conclusion, we have presented a systematic view of acceleration methods for GNNs on large-scale datasets. Our survey covers GNN training and inference algorithms, efficient GNN systems, and customized hardware for GNNs. Additionally, we have discussed the special cases when GNNs are applied to heterogeneous graphs and dynamic graphs. Acceleration methods pave the way for more efficient GNN pipelines, enabling the use of larger GNN models for solving more complex tasks on graphs.
Acknowledgements.
We thank Linghao Song for the helpful discussions and comments throughout this work. We thank Xiang Song and Kezhao Huang for providing helpful feedback on an early draft of the paper. This work is partially supported by NSF (2211557, 1937599, 2119643), NASA, SRC, Okawa Foundation Grant, Amazon Research Awards, Cisco Research Grant, Picsart Gifts, and Snapchat Gifts.